### v1

# Matching-Based Policy Learning

Abstract
Treatment heterogeneity is ubiquitous in many areas, motivating practitioners to search for the optimal policy that maximizes the expected outcome based on individualized characteristics. However, most existing policy learning methods rely on weighting-based approaches, which may suffer from high instability in observational studies. To enhance the robustness of the estimated policy, we propose a matching-based estimator of the policy improvement upon a randomized baseline. After correcting the conditional bias, we learn the optimal policy by maximizing the estimate over a policy class. We derive a non-asymptotic high probability bound for the regret of the learned policy and show that the convergence rate is almost $1/\sqrt{n}$. The competitive finite sample performance of the proposed method is demonstrated in extensive simulation studies and a real data application.
Keywords: individualized treatment rules, matching, observational study, policy learning, regret bound
1 Introduction
Recently, data-driven personalized decision-making has received increased attention across various fields. In healthcare, clinicians utilize patients’ characteristics and medical history to tailor diabetes management, aiming to enhance health outcomes (Bertsimas et al., 2017). In public policy, leveraging individualized features can lead to an improved allocation of social services and thus enhance overall efficiency (Kube et al., 2019). In product recommendations, the online platforms can effectively keep a significant number of customers engaged by recommending the products that match their preferences (Bastani et al., 2022). A common feature underlying these applications is the presence of treatment heterogeneity, which is a blessing that brings opportunities to improve the expected outcome in the target population. Therefore, it is imperative to exploit such heterogeneity efficiently and learn the optimal policy that allocates treatments based on individualized characteristics.
Statistical methods for estimating the optimal policy have been extensively developed in the literature, among which can be roughly divided into model-based methods and direct-search methods (e.g., Qian and Murphy, 2011; Zhao et al., 2012; Zhou et al., 2017; Kitagawa and Tetenov, 2018; Shi et al., 2018). For instance, Q-learning (Qian and Murphy, 2011) and A-learning (Shi et al., 2018) first fit the mean outcome given covariates and treatment, or posit a model for the contrast function, then determine the optimal policy based on the corresponding estimate. Both methods belong to model-based approaches, and their effectiveness relies heavily on correct model specification. Alternatively, outcome weighted learning (O-learning, Zhao et al., 2012) is one of the pioneer works of direct-search methods that nonparametrically learns the optimal policy by utilizing the inverse probability weighting estimator (IPWE) for the mean outcome corresponding to the policy, i.e., value function. In observational studies, the unknown propensity score in O-learning needs to be estimated using logistic regression or machine learning approaches (Lee et al., 2010) for adjusting confounding bias. One recognized potential drawback of weighting-based approaches is the instability caused by extreme estimated propensity scores or model misspecification.
To provide additional robustness, some policy learning literature based on augmented IPWE (AIPWE) of value function has emerged (e.g., Zhang et al., 2012a, b; Zhao et al., 2019; Athey and Wager, 2021; Pan and Zhao, 2021), which combines the strengths of both the aforementioned strategies. Zhang et al. (2012b) proposed an AIPWE of value function that possesses doubly robustness in the sense that the estimator is consistent if either propensity score model or outcome mean model is correctly posited, where parametric models were used for these nuisance functions. Yet, the performance could be unstable when both models are misspecified (Kang and Schafer, 2007). The seminal work by Athey and Wager (2021) leverages cross-fitting technique (Chernozhukov et al., 2018) and utilizes flexible machine learning approaches for nuisance functions. They demonstrated that the learned policy achieves minimax optimal regret if the mean-squared errors of nuisance functions satisfy some order conditions of convergence rate. Though the promising theoretical guarantee, this AIPWE-based method still suffers from extreme weights, especially when the sample size is relatively small. Similar phenomena have been discussed in both policy learning and causal inference literature (e.g., Zhao et al., 2019; Wu et al., 2022). Hence, it is crucial to develop a robust approach for policy learning in observational studies with favorable finite sample performance.
In this paper, we propose a matching-based method (MB-learning) for learning the optimal policy with observational data. Matching is an appealing approach for estimating causal effect in observational studies, which balances covariates between different groups intuitively and exhibits robustness against extreme weights. There is a significant amount of work on matching estimators (see, e.g., Imbens and Rubin (2015) and Rosenbaum (2020) for literature review). The insightful idea behind MB-learning originates from an alternative form of the value function based on the pair of potential outcomes (referred to as advantage function in Section 2), which can be interpreted as the value function improvement upon a randomized baseline. Motivated by the advantage function, we impute the missing counterfactual outcomes to develop the corresponding estimator, where we utilize nearest-neighbor (NN) matching on covariates (Abadie and Imbens, 2006). One recognized limitation of matching estimator is its susceptibility to the curse of dimensionality, as it is not $\sqrt{n}$-consistent in general due to the conditional bias. To mitigate this issue, we employ a bias correction technique proposed by Abadie and Imbens (2011) and develop an improved matching-based estimator with $\sqrt{n}$-asymptotic normality. Leveraging the enhanced estimator, we recast the policy learning optimization into a weighted classification problem and derive the optimal policy.
This paper makes four major contributions to existing literature:
  1. 1.
    Distinct from the (A)IPWE-based approaches, MB-learning is a robust method for learning the optimal policy in observational studies. By utilizing a bias-corrected matching-based estimator, MB-learning mitigates the potential variability arising from propensity score model misspecification or extreme weights.
  2. 2.
    We establish theoretical guarantee for the optimal policy learned by the matching-based estimator, where a non-asymptotic high probability bound for the value function difference between the best-in-class policy and the learned policy is derived.
  3. 3.
    We demonstrate the superiority of MB-learning both theoretically and empirically. Specifically, Athey and Wager (2021) established that their approach achieves rate-optimality in terms of minimax regret, which is on the order of $1/\sqrt{n}$. In contrast, the regret bound of MB-learning is almost $1/\sqrt{n}$. However, as illustrated in simulation studies and real data analysis, MB-learning exhibits more robust finite sample performance.
  4. 4.
    To our knowledge, matching has rarely been utilized for policy learning. The only exception is matched learning (Wu et al., 2020, M-learning), where they estimated the optimal policy by weighted support vector machine and presented the competitive performance in observational studies. We highlight several advantages of MB-learning. First, MB-learning is more intuitive that directly targets an estimator of the advantage function, while M-learning seeks to maximize an empirical analogue of a complicated objective function (Wu et al., 2020, Theorem 3.1). Second, the policy derived from MB-learning possesses appealing interpretability and implementability for practitioners, while that from M-learning with Gaussian kernel lacks of interpretation. Third, the presence of bias introduced by NN matching could substantially deteriorate the performance of M-learning when dealing with moderately high-dimensional covariates.
The rest of this paper is organized as follows. In Section 2, we develop the methodology of MB-learning. Specifically, we formulate the treatment assignment framework and introduce the AIPWE-based policy learning approach developed by Athey and Wager (2021). Then we propose a matching-based estimator of the advantage function, which is further improved by a bias correction technique. Finally, the optimal policy is derived from the resulting estimate. In Section 3, we establish rigorous theoretical guarantee for MB-learning. The empirical performance of MB-learning is demonstrated through extensive simulation studies in Section 4, while the application to National Supported Work (NSW) Program is presented in Section 5. We conclude this paper with discussion in Section 6.
2 Methodology
2.1 Notations and Preliminaries
[…]
We consider the problem of learning a policy $\pi$, a decision function from covariates space $\mathcal{X}$ to treatment space $\{0,1\}$, with observational data. In practice, it is unrealistic to assign a complex policy to the target population. We should attach problem-specific constraints to our learned policy, such as political restrictions (Kitagawa and Tetenov, 2018), fairness (Athey and Wager, 2021) and interpretability (Zhang et al., 2012a). Therefore, we seek to learn the optimal policy within a restricted policy class $\Pi$. The value function $V(\pi)$ relative to $\pi$ (Zhao et al., 2012) is the expected outcome after implementing the treatment allocation by $\pi$, defined as $V(\pi)=E\left[Y(\pi(X))\right]$. The optimal policy $\pi^{*}$ within class $\Pi$ maximizes the value function $V(\pi)$, i.e., $\pi^{*}=\arg\max_{\pi\in\Pi}V(\pi)$.
The advantage function serves as an alternative form of the value function:
$A(\pi)=E\left[\left(2\pi(X)-1\right)\tau(X)\right].$
Since $V(\pi)=E[Y(0)+(Y(1)-Y(0))\pi(X)]$, it follows that $A(\pi)=2\left[V(\pi)-\frac{E[Y(1)]+E[Y(0)]}{2}\right]$ by algebraic manipulations. Hence, the advantage function $A(\pi)$ can be interpreted as the improvement achieved by implementing policy $\pi$ relative to a randomly assigned policy with equal probability (Mo et al., 2021). Moreover, the policy that maximizes $A(\pi)$ is equivalent to that maximizes $V(\pi)$. Throughout this paper, we focus on policy learning based on $A(\pi)$.
Following the idea of semiparametric efficient estimation (Robins et al., 1994), some existing literature developed AIPWEs of $A(\pi)$ and subsequently determined the optimal policies based on these estimators (e.g., Zhang et al., 2012a; Athey and Wager, 2021). Notably, the AIPWE proposed by Athey and Wager (2021) is $\hat{A}_{aipw}(\pi)=n^{-1}\sum_{i=1}^{n}\left(2\pi(X_{i})-1\right)\Gamma_{i}$, where
$\Gamma_{i}=\mu(X_{i},1)-\mu(X_{i},0)+\frac{W_{i}-e(X_{i})}{e(X_{i})\left(1-e(X_{i})\right)}\left(Y_{i}-\mu(X_{i},W_{i})\right).$
The advantage of this AIPWE for $A(\pi)$ is its doubly robustness, providing additional protection from model misspecification in observational studies. Athey and Wager (2021) employed black-box machine learning methods, which achieve sufficiently fast convergence rate in mean-squared error, to estimate the nuisance functions. They developed a plug-in estimator of $\hat{A}_{aipw}(\pi)$ and learned the optimal policy based on the resulting estimator. However, as previously mentioned, estimating the unknown propensity score $e(x)$ in the denominator of $\Gamma_{i}$ may introduce instability, particularly when the estimated propensity score is extreme. In addition, the estimation for nuisance functions by machine learning methods may be unsatisfactory with a relatively small sample size, even though the performance can be guaranteed in large sample. Therefore, it is imperative to develop a more sensible approach for learning the optimal policy in observational studies, while ensuring desirable efficiency and robustness in finite sample.
2.2 Matching-Based Advantage Function
Our proposed approach is inspired by an alternative expression of the advantage function:
$A(\pi)=E\left[\left(2\pi(X)-1\right)\left(Y(1)-Y(0)\right)\right]=E\left[\left(2\pi(X)-1\right)\operatorname{sign}\left(Y(1)-Y(0)\right)\left|Y(1)-Y(0)\right|\right].$
The second equality indicates that, in order to maximize $A(\pi)$, one should assign each unit to the treatment group in which its potential outcome can be larger; meanwhile, a more significant discrepancy between the pair of potential outcomes contributes to a greater influence in the treatment allocation. Building on this intuition, we seek to develop an estimator of $A(\pi)$ based on the pair of potential outcomes. However, since only one of the pair of potential outcomes is observable, we utilize matching estimator to impute the missing counterfactual outcome.
We leverage the matching estimator proposed by Abadie and Imbens (2006), where they employed matching with replacement, allowing each unit to be used as a match multiple times. Specifically, for any $x\in\mathcal{X}$ and some positive definite matrix $V$, let $\|x\|=(x^{\top}Vx)^{1/2}$ denote a vector norm. We define the matched set $\mathcal{J}_{M}(i)$ as the set that contains the indices of the first $M$ matches for unit $i$:
$\mathcal{J}_{M}(i)=\left\{j:W_{j}=1-W_{i},\sum_{k:W_{k}=1-W_{i}}I\{\|X_{k}-X_{i}\|\leq\|X_{j}-X_{i}\|\}\leq M\right\}.$
Here, we ignore the ties in matching, which happens with probability zero if $X$ is continuously distributed (Abadie and Imbens, 2006). In this paper, we consider the Mahalanobis metric $V_{\text{maha}}$, which is commonly used in practice:
$V_{\text{maha}}=\left(\frac{1}{n}\sum_{i=1}^{n}(X_{i}-\bar{X})(X_{i}-\bar{X})^{\top}\right)^{-1},\text{ where }\bar{X}=\frac{1}{n}\sum_{i=1}^{n}X_{i}.$
The Mahalanobis metric is not the only choice, see Abadie and Imbens (2006) for discussion on other metrics. The matching estimator of potential outcomes is given by
$\hat{Y}_{i}(0)=\left\{\begin{aligned} &Y_{i}&\text{ if }W_{i}=0,\\ &\frac{1}{M}\underset{j\in\mathcal{J}_{M}(i)}{\sum}Y_{j}&\text{ if }W_{i}=1,\end{aligned}\right.\quad\text{ and }\quad\hat{Y}_{i}(1)=\left\{\begin{aligned} &\frac{1}{M}\underset{j\in\mathcal{J}_{M}(i)}{\sum}Y_{j}&\text{ if }W_{i}=0,\\ &Y_{i}&\text{ if }W_{i}=1.\end{aligned}\right.$
For each unit $i$, the pair of matching estimators $\left(\hat{Y}_{i}(1),\hat{Y}_{i}(0)\right)$ mimics the pair of potential outcomes $\left(Y_{i}(1),Y_{i}(0)\right)$. Finally, we propose the matching-based advantage function:
$\hat{A}_{match}(\pi)=\frac{1}{n}\sum^{n}_{i=1}\left(2\pi(X_{i})-1\right)\left(\hat{Y}_{i}(1)-\hat{Y}_{i}(0)\right).$ (1)
In the following, we will establish the consistency and asymptotic normality of the matching-based estimator $\hat{A}_{match}(\pi)$, with subsequent discussion of its limitation and the proposal for improvement. Let $M_{i}=(M_{i1},\ldots,M_{in})^{\top}$ denote the $n$-dimensional vector indicating the units in $\mathcal{J}_{M}(i)$, where $M_{ij}=1$ if unit $j\in\mathcal{J}_{M}(i)$, and $M_{ij}=0$ otherwise, i.e., $M_{ij}=I\left\{j\in\mathcal{J}_{M}(i)\right\}$. For convenience of further analysis, we formulate $\hat{A}_{match}(\pi)$ as a linear combination of $\{Y_{i}\}_{i=1}^{n}$ after basic manipulations.
Lemma 1.
The matching-based advantage function $\hat{A}_{match}(\pi)$ has the following expression:
$\hat{A}_{match}(\pi)=\frac{1}{n}\sum^{n}_{i=1}(2W_{i}-1)\left[(2\pi(X_{i})-1)+\frac{K_{M}(\pi,i)}{M}\right]Y_{i},$
where $K_{M}(\pi,i)=\sum^{n}_{j=1}\left(2\pi(X_{j})-1\right)M_{ji}=\sum_{j:M_{ji}=1}\left(2\pi(X_{j})-1\right)$.
It is noteworthy that $\{K_{M}(\pi,i)\}_{i=1}^{n}$ is the key component for asymptotic analysis. The properties of $\hat{A}_{match}(\pi)$ resemble those of the matching estimator for $\tau$ proposed by Abadie and Imbens (2006):
$\hat{\tau}=\frac{1}{n}\sum^{n}_{i=1}\left(\hat{Y}_{i}(1)-\hat{Y}_{i}(0)\right)=\frac{1}{n}\sum^{n}_{i=1}(2W_{i}-1)\left(1+\frac{K_{M}(i)}{M}\right)Y_{i},$
where $K_{M}(i)=\sum_{j=1}^{n}M_{ji}$ is the number of times unit $i$ has been matched. When $\pi(x)\equiv 1$, it follows that $\hat{A}_{match}(\pi)=\hat{\tau}$ and $K_{M}(\pi,i)=K_{M}(i)$. Therefore, the matching estimator $\hat{\tau}$ is a special case of our proposed estimator. Furthermore, we proceed to show that $K_{M}(\pi,i)$ has the same order as $K_{M}(i)$ under the following assumption.
Assumption 1.
The $p$-dimensional vector $X$ is continuously distributed on a convex and compact support $\mathcal{X}_{0}\subset\mathbb{R}^{p}$, and its density is bounded and bound away from zero on $\mathcal{X}_{0}$.
Assumption 1 is a standard regularity condition for the matching estimator of average causal effect (Abadie and Imbens, 2006), supposing $X$ is continuously distributed. In contrast, discrete covariates with finite support points can be addressed by analyzing the estimation of the advantage function within subsamples. However, for high-dimensional and continuous $X$, the assumption that its density is bound away from zero may be too stringent (Otsu and Rai, 2017).
Lemma 2.
Under Assumption 1, for any $\pi\in\Pi$, $K_{M}(\pi,i)=O_{p}(1)$ and $E[K_{M}(\pi,i)^{q}]$ is bounded uniformly in $n$ for any $q>0$.
Analyzing the asymptotic properties of $\hat{A}_{match}(\pi)$ may be challenging due to the dependence among $\{K_{M}(\pi,i)\}^{n}_{i=1}$. Inspired by Abadie and Imbens (2006), we decompose $\hat{A}_{match}(\pi)$ into several components for further analysis.
Lemma 3.
The matching-based advantage function $\hat{A}_{match}(\pi)$ has the following decomposition:
$\hat{A}_{match}(\pi)=\bar{A}(\pi)+E_{M}(\pi)+B_{M}(\pi),$
where
$\bar{A}(\pi)=\frac{1}{n}\sum^{n}_{i=1}(2\pi(X_{i})-1)(\mu(X_{i},1)-\mu(X_{i},0)),$
$E_{M}(\pi)=\frac{1}{n}\sum^{n}_{i=1}(2W_{i}-1)\left[(2\pi(X_{i})-1)+\frac{K_{M}(\pi,i)}{M}\right]\epsilon_{i},$
$B_{M}(\pi)=\frac{1}{n}\sum^{n}_{i=1}(2W_{i}-1)(2\pi(X_{i})-1)\frac{1}{M}\underset{j\in\mathcal{J}_{M}(i)}{\sum}(\mu(X_{i},1-W_{i})-\mu(X_{j},1-W_{i})).$ (2)
Let $\mathbf{X}$ denote the $n\times p$ matrix whose $i$th row is $X^{\top}_{i}$, and $\mathbf{W}$ is the $n\times 1$ vector with the $i$th element equal to $W_{i}$. It follows that $B_{M}(\pi)=E\left[\hat{A}_{match}(\pi)-\bar{A}(\pi)\mid\mathbf{X},\mathbf{W}\right]$ after some algebra, referred to as the conditional bias relative to $\bar{A}(\pi)$. Utilizing the decomposition in Lemma 3, we proceed to derive the large sample properties of $\hat{A}_{match}(\pi)$ by analyzing its components separately. First, we derive the order of $B_{M}(\pi)$ in the following lemma.
Lemma 4.
Under Assumption 1, suppose that $\mu(x,1)$ and $\mu(x,0)$ are Lipschitz in $\mathcal{X}_{0}$, then for any $\pi\in\Pi$, $B_{M}(\pi)=O_{p}\left(n^{-1/p}\right)$.
Lemma 4 demonstrates that, although the conditional bias is asymptotically negligible, its convergence rate can be slow with a moderately large $p$. Let $V^{\tau(X)}(\pi)=E\left[(2\pi(X)-1)\tau(X)-A(\pi)\right]^{2}$. The conditional variance of $E_{M}(\pi)$ given $\mathbf{X}$ and $\mathbf{W}$ is
$\text{Var}\left[E_{M}(\pi)\mid\mathbf{X},\mathbf{W}\right]=\frac{1}{n^{2}}\sum^{n}_{i=1}\left[(2\pi(X_{i})-1)+\frac{K_{M}(\pi,i)}{M}\right]^{2}\sigma^{2}(X_{i},W_{i}),$
and the corresponding normalized conditional variance is denoted by $V^{E}(\pi)=n\text{Var}\left[E_{M}(\pi)\mid\mathbf{X},\mathbf{W}\right]$. Under the following regularity conditions, we establish the consistency and asymptotic normality of $\hat{A}_{match}(\pi)$.
Assumption 2.
Assume the following conditions:
(i) For $w=0,1$, $\mu(x,w)$ and $\sigma^{2}(x,w)$ are Lipschitz in $\mathcal{X}_{0}$.
(ii) For $w=0,1$, E$\left[Y^{4}\mid X=x,W=w\right]$ is bounded uniformly in $\mathcal{X}_{0}$.
(iii) For any $\pi\in\Pi$, $V^{E}(\pi)$ bounded away from zero in the Cartesian product space $\mathcal{X}_{0}^{n}\times\{0,1\}^{n}$.
Assumption 2(i) is a mild condition for restricting the smoothness and boundness of conditional mean and variance. Assumption 2(ii) is a moment condition for the central limit theorem. Assumption 2(iii) is analogous to the Assumption 4(iii) in Abadie and Imbens (2006), which guarantees the nonsingularity of $V^{E}(\pi)$ uniformly in $\Pi$.
Theorem 1.
(i) Under Assumptions 1 and 2(i), then for any $\pi\in\Pi$,
$\hat{A}_{match}(\pi)-A(\pi)\stackrel{{\scriptstyle p}}{{\rightarrow}}0.$
(ii) Under Assumptions 1 and 2, then for any $\pi\in\Pi$,
$\left(V^{\tau(X)}(\pi)+V^{E}(\pi)\right)^{-1/2}\sqrt{n}\left(\hat{A}_{match}(\pi)-B_{M}(\pi)-A(\pi)\right)\stackrel{{\scriptstyle d}}{{\rightarrow}}N(0,1).$
Theorem 1 shows that the matching-based estimator $\hat{A}_{match}(\pi)$ converges in probability to the advantage function $A(\pi)$ regardless of covariates dimension $p$. In addition, $\hat{A}_{match}(\pi)$ is $\sqrt{n}$-asymptotically normal distributed after subtracting the conditional bias $B_{M}(\pi)$. Furthermore, leveraging Lemma 4 and Theorem 1, we establish that
$\hat{A}_{match}(\pi)-A(\pi)=O_{p}\left(n^{-1/\max\{2,p\}}\right).$
When the covariate is scalar, $\hat{A}_{match}(\pi)$ possesses $\sqrt{n}$-asymptotic normality. However, the convergence rate can be slow with a moderately large $p$, where $B_{M}(\pi)$ becomes the leading term. Therefore, it is imperative to improve the estimator $\hat{A}_{match}(\pi)$ that eliminates the conditional bias and thus remedies the slow rate.
2.3 Bias-Corrected Matching-Based Advantage Function
To alleviate the conditional bias in $\hat{\tau}$, Abadie and Imbens (2011) utilized a power series estimator of the outcome mean model $\mu(x,w)$ to develop a bias-corrected matching estimator. Specifically, let $\lambda=\left(\lambda_{1},\ldots,\lambda_{p}\right)^{\top}$ denote a $p$-dimensional vector of nonnegative integers with the norm $\|\lambda\|_{1}=\sum_{j=1}^{p}\lambda_{j}$. Define that $x^{\lambda}=\prod_{j=1}^{p}(x_{j})^{\lambda_{j}}$, where $x_{j}$ is the $j$th element of $p$-dimensional vector $x$. For the series $\{\lambda(j)\}_{j=1}^{\infty}$ containing all distinct vectors with nondecreasing $\|\lambda(j)\|_{1}$, define the $K$-dimensional vector $p^{K}(x)=\left(x^{\lambda(1)},\ldots,x^{\lambda(K)}\right)^{\top}$. The power series estimator (Newey, 1997) of $\mu(x,w)$ is given by
$\hat{\mu}(x,w)=\left(p^{K}(x)\right)^{\top}\left(\underset{i:W_{i}=w}{\sum}p^{K}(X_{i})\left(p^{K}(X_{i})\right)^{\top}\right)^{-}\underset{i:W_{i}=w}{\sum}p^{K}(X_{i})Y_{i},$
where $(\cdot)^{-}$ is a generalized inverse. Finally, the bias-corrected matching estimator of potential outcomes is defined as
$\tilde{Y}_{i}(0)=\left\{\begin{aligned} &Y_{i}&\text{ if }W_{i}=0,\\ &\frac{1}{M}\underset{j\in\mathcal{J}_{M}(i)}{\sum}\left(Y_{j}+\hat{\mu}(X_{i},0)-\hat{\mu}(X_{j},0)\right)&\text{ if }W_{i}=1,\end{aligned}\right.$
and
$\tilde{Y}_{i}(1)=\left\{\begin{aligned} &\frac{1}{M}\underset{j\in\mathcal{J}_{M}(i)}{\sum}\left(Y_{j}+\hat{\mu}(X_{i},1)-\hat{\mu}(X_{j},1)\right)&\text{ if }W_{i}=0,\\ &Y_{i}&\text{ if }W_{i}=1.\end{aligned}\right.$
Remarkably, they established the $\sqrt{n}$-asymptotic normality of the corresponding matching estimator $\hat{\tau}_{bc}=n^{-1}\sum_{i=1}^{n}\left(\tilde{Y}_{i}(1)-\tilde{Y}_{i}(1)\right)$ for $\tau$ under regularity conditions. Motivated by Abadie and Imbens (2011), we develop an improved matching-based advantage function:
$\hat{A}^{bc}_{match}(\pi)=\frac{1}{n}\sum^{n}_{i=1}\left(2\pi(X_{i})-1\right)\left(\tilde{Y}_{i}(1)-\tilde{Y}_{i}(0)\right).$ (3)
After basic manipulations, it follows that $\hat{A}^{bc}_{match}(\pi)=\hat{A}_{match}(\pi)-\hat{B}_{M}(\pi)$, where $\hat{B}_{M}(\pi)$ serves as an estimator of $B_{M}(\pi)$ given by
$\hat{B}_{M}(\pi)=\frac{1}{n}\sum^{n}_{i=1}(2W_{i}-1)\left(2\pi(X_{i})-1\right)\frac{1}{M}\underset{j\in\mathcal{J}_{M}(i)}{\sum}\left(\hat{\mu}(X_{i},1-W_{i})-\hat{\mu}(X_{j},1-W_{i})\right).$
We proceed to establish the $\sqrt{n}$-asymptotic normality of $\hat{A}^{bc}_{match}(\pi)$ in Corollary 2, which implies that $\hat{A}^{bc}_{match}(\pi)-A(\pi)=O_{p}(1/\sqrt{n})$. Assumption 3 contains regularity conditions for the power series estimator $\hat{\mu}(x,w)$ to ensure its fast convergence rate and thus eliminate the conditional bias. Detailed discussions can be found in Abadie and Imbens (2011) and Otsu and Rai (2017).
Assumption 3.
Assume the following conditions:
(i) The support $\mathcal{X}_{0}$ is a Cartesian product of compact intervals.
(ii) $K=O(n^{v})$, where $0<v<\min\{2/(4p+3),2/(4p^{2}-p)\}$.
(iii) For each $\lambda$ and $w=0,1$, the derivative $\partial^{\|\lambda\|_{1}}\mu(x,w)/\partial x^{\lambda_{1}}_{1}\cdots\partial x^{\lambda_{p}}_{p}$ exists and its norm is bounded by $C^{\|\lambda\|_{1}}$ for some positive constant $C$.
Corollary 2.
Under Assumptions 1-3, then for any $\pi\in\Pi$,
$\left(V^{\tau(X)}(\pi)+V^{E}(\pi)\right)^{-1/2}\sqrt{n}\left(\hat{A}^{bc}_{match}(\pi)-A(\pi)\right)\stackrel{{\scriptstyle d}}{{\rightarrow}}N(0,1).$
2.4 From Estimating to Learning
In the preceding discussion, we have introduced a bias-corrected matching-based estimator $\hat{A}^{bc}_{match}(\pi)$ of the advantage function $A(\pi)$ with a fixed policy $\pi$. We proceed to demonstrate the procedure of determining the optimal policy based on $\hat{A}^{bc}_{match}(\pi)$. Notice that $\hat{A}^{bc}_{match}(\pi)$ has the following expression:
$\hat{A}^{bc}_{match}(\pi)=\frac{1}{n}\sum^{n}_{i=1}\left(2\pi(X_{i})-1\right)\operatorname{sign}\left(\tilde{Y}_{i}(1)-\tilde{Y}_{i}(0)\right)\left|\tilde{Y}_{i}(1)-\tilde{Y}_{i}(0)\right|.$ (4)
Therefore, the maximization for learning the optimal policy is equivalent to a weighted classification problem, where we can treat $\pi(\cdot)$ as the classifier to be trained, $\left\{\operatorname{sign}\left(\tilde{Y}_{i}(1)-\tilde{Y}_{i}(0)\right)\right\}^{n}_{i=1}$ as the labels, and $\left\{\left|\tilde{Y}_{i}(1)-\tilde{Y}_{i}(0)\right|\right\}^{n}_{i=1}$ as the weights.
Given a set of $n$ training samples $\{X_{i},W_{i},Y_{i}\}_{i=1}^{n}$ and a restricted policy class $\Pi$, we summarize the implementation of MB-learning as follows:
  • Impute the pairs of potential outcomes $\left\{Y_{i}(1),Y_{i}(0)\right\}^{n}_{i=1}$ using the bias-corrected matching estimators $\left\{\tilde{Y}_{i}(1),\tilde{Y}_{i}(0)\right\}^{n}_{i=1}$ defined in Section 2.3.
  • Form the bias-corrected matching-based advantage function $\hat{A}^{bc}_{match}(\pi)$.
  • Implement the weighted classification (4) to learn the optimal policy $\hat{\pi}=\arg\max_{\pi\in\Pi}\hat{A}^{bc}_{match}(\pi)$.
The idea of recasting the maximization of value function into a weighted classification framework has been previously discussed (e.g., Zhang et al., 2012a; Zhao et al., 2012; Athey and Wager, 2021). However, solving this maximization may be problematic since it is not a convex optimization problem. Following the argument in Athey and Wager (2021), we set the policy class $\Pi$ as a class of fixed-depth decision trees and implement the weighted classification problem using R package policytree (Sverdrup et al., 2020; R Core Team, 2023). This approach offers appealing interpretability and practical implementation for decision-makers. Some existing literature has also considered using convex surrogate loss function to replace the nonsmooth indicator function, thereby reformulating the policy learning problem into a convex optimization problem and implementing the optimization using off-the-shelf packages (Zhao et al., 2012; Liu et al., 2018).
3 Regret Bound
In the previous section, we have investigated asymptotic behaviour of the proposed matching-based estimators for the advantage function $A(\pi)$ with a fixed $\pi$, focusing on the estimation properties. However, in statistical learning literature, one commonly used criterion to evaluate a policy is how effectively it minimizes the gap between the value function of the optimal policy and the policy we have learned (e.g., Zhao et al., 2012; Kitagawa and Tetenov, 2018; Athey and Wager, 2021; Zhou et al., 2023). This criterion is formalized through the concept of regret. Specifically, the utilitarian regret of a policy $\pi\in\Pi$ (Athey and Wager, 2021; Zhou et al., 2023) is defined as
$R(\pi)=V(\pi^{*})-V(\pi).$
We proceed to investigate the convergence rate of $R(\hat{\pi})$, i.e., we will derive a high probability upper bound for the regret. Let $\text{VC}(\Pi)$ denote the Vapnik–Chervonenkis (VC) dimension (van der Vaart and Wellner, 2023) of the policy class $\Pi$. We assume that $\text{VC}(\Pi)$ is finite in Theorem 3 to control the complexity of $\Pi$. Suppose that we learn the optimal policy within a class of depth-$H$ decision trees in the implementation of MB-learning, then the corresponding VC dimension is $\tilde{O}\left(2^{H}\log p\right)$ (Athey and Wager, 2021). Here, the notation $\tilde{O}\left(f(n)\right)$ represents any factor bounded above by $g(f(n))f(n)$, where the function $g(\cdot)$ scales polylogarithmically with its argument. We also assume the countability of $\Pi$, which is a standard technical condition to simplify the measurability in the proof (e.g., Kitagawa and Tetenov, 2018). Let the notation $A\lesssim B$ denote that $A$ is less than $B$ up to a constant.
Theorem 3.
Under Assumptions 1 and 3, suppose that the outcome $|Y|\leq L$ almost surely for some positive constant $L$, and the policy class $\Pi$ is countable with finite $\text{VC}(\Pi)$, then for any $\delta>0$ and any $q>1$, there exists an integer $N_{\delta,q}$ such that with at least probability $1-11\delta$ and for all $n>N_{\delta,q}$,
$R(\hat{\pi})\lesssim\frac{L}{\sqrt{Mn^{1-1/q}}}\left(\sqrt{\text{VC}(\Pi)}+\sqrt{\frac{\log\frac{1}{\delta}}{M}}\right)+\frac{1}{\sqrt{n}}\left(\sqrt{\text{VC}(\Pi)}+\sqrt{\log\frac{1}{\delta}}\right)\left(\sqrt{V^{*}}+L\right)+\frac{L\log\frac{1}{\delta}}{Mn^{1-1/q}},$
where $V^{*}=4\sup_{\pi_{1},\pi_{2}\in\Pi}E\left[\left(\mu(X,1)-\mu(X,0)\right)^{2}\left(\pi_{1}(X)-\pi_{2}(X)\right)^{2}\right].$
It suffices to analyze the convergence rate of $A(\pi^{*})-A(\hat{\pi})$ in the proof, where we decompose it into several terms and derive the high probability upper bounds separately. We note that it could be challenging to derive the upper bound for
$E\left[\underset{\pi\in\Pi}{\sup}\left|\frac{1}{n}\sum^{n}_{i=1}(2W_{i}-1)\left[\sum_{j=1}^{n}\pi(X_{j})M_{ji}\right]\epsilon_{i}\right|\right]$ (5)
due to the dependence among the terms $\left\{\sum_{j=1}^{n}\pi(X_{j})M_{ji}\right\}_{i=1}^{n}$, which is originated from the use of matching estimator. However, conditional on $\{X_{i},W_{i},\epsilon_{i}\}_{i=1}^{n}$, we can still utilize standard empirical process techniques including symmetrization (Boucheron et al., 2013) and demonstrate that the corresponding Rademacher process is sub-Gaussian. Consequently, we can leverage Dudley’s entropy integral bound (Vershynin, 2018) and the order of $\max_{i\leq n}K_{M}(i)$ to derive an upper bound for (5).
Theorem 3 shows that the regret bound depends on the VC dimension $\text{VC}(\Pi)$, the number of matches $M$, and the quantity $V^{*}$. A better regret bound can be achieved if we decrease the complexity of the policy class or increase the size of matching. Analogous phenomenon concerning the relationship between the regret $R(\hat{\pi})$ and $M$ has been observed in Abadie and Imbens (2006, Theorem 5), where they stated that a larger $M$ can lead to a more efficient matching estimator of $\tau$. Recall that the unit-level term of $\bar{A}(\pi)$ is given by $\bar{A}_{i}(\pi)=(2\pi(X_{i})-1)\left(\mu(X_{i},1)-\mu(X_{i},0)\right)$, then $V^{*}$ can be reformulated as $\sup_{\pi_{1},\pi_{2}\in\Pi}E\left[\bar{A}_{i}(\pi_{1})-\bar{A}_{i}(\pi_{2})\right]^{2}$. Therefore, we can interpret $V^{*}$ as the worst-case discrepancy between the advantage functions of any two policies in the class $\Pi$. See Zhou et al. (2023) for detailed discussion.
From Theorem 3 we establish that the convergence rate of regret is
$R(\hat{\pi})=O_{p}\left(\sqrt{\frac{\text{VC}(\Pi)}{n^{1-1/q}}}\right),$
where $q>1$. We let $q$ go to $\infty$, then the convergence rate almost achieves $1/\sqrt{n}$. Though the convergence rate of MB-learning is slower than that of the method proposed by Athey and Wager (2021), which is exactly $1/\sqrt{n}$, our results in simulation studies and real data analysis demonstrate the superiority of the proposed method, especially in a relatively small sample size. Recall that M-learning proposed by Wu et al. (2020) also utilizes the matching for learning the optimal policy. However, they established the convergence rate based on the results of NN estimator (e.g., Devroye et al., 2013), where the number of matches grows with sample size. In contrast, the convergence rate of MB-learning is developed with a fixed number $M$. Additionally, Kitagawa and Tetenov (2018) showed that their IPWE-based method has a regret bound faster than $1/\sqrt{n}$ under the margin assumption, which is commonly used in classification literature (e.g., Tsybakov, 2004; Massart and Nédélec, 2006). Therefore, a faster convergence rate of MB-learning may be achieved under this assumption, but it is beyond the scope of this work.
4 Simulation Studies
4.1 Data Generating Process
We conduct extensive simulation studies to assess the finite sample performance of MB-learning. In each simulation setup, the covariates vector $X=(X_{1},X_{2},X_{3},X_{4})^{\top}$ is independently generated from $N(0,\mathbf{I}_{4})$. The binary treatment $W$ is generated from $\Pr(W=1|X)=\exp\left(l(X)\right)/\left[1+\exp\left(l(X)\right)\right]$, where we consider five scenarios of propensity score models based on different specifications of $l(X)$. Detailed information is presented in Table S1 of Supplementary Materials. The outcome variable is generated by $Y=m(X)+Wc(X)+Z,$ where $Z$ is the noisy variable independently generated from $N(0,1)$. We consider two scenarios of the main effect model $m(\cdot)$: Linear, $m(X)=1+2X_{1}-X_{2}+0.5X_{3}-1.5X_{4}$, and Nonlinear, $m(X)=4\sin(X_{1})+2.5\cos(X_{2})-X_{3}X_{4}$. The first main effect model is linear with its argument, while the second model exhibits highly nonlinearity. Additionally, two scenarios of the contrast function $c(\cdot)$ are also considered: Tree, $c(X)=2I\left\{X_{1}>0,X_{2}>0\right\}-1$, and Non-tree, $c(X)=2I\left\{2X_{2}-\exp(1+X_{1})+2>0\right\}-1$. We define the global optimal policy $\pi^{*}_{opt}$ as the optimal policy within all measurable functions from $\mathcal{X}$ to $\{0,1\}$ that maximizes the value function (Mo et al., 2021), i.e., $\pi^{*}_{opt}=\arg\max_{\pi}V(\pi)$. Note that the contrast function $c(X)$ determines the structure of the global optimal policy since $\pi^{*}_{opt}(X)=I\left\{c(X)>0\right\}$. Therefore, in the first scenario, we proceed to assess the performance of MB-learning when the restricted policy class $\Pi$ includes the global optimal policy $\pi^{*}_{opt}$. On the other hand, we investigate how the misspecification of $\Pi$ influences the proposed method in the second scenario. Similar simulation studies have also been conducted in the literature concerning tree-based policy learning approaches (e.g., Zhang et al., 2012a; Laber and Zhao, 2015; Athey and Wager, 2021).
4.2 Compared Methods and Implementation Details
The proposed method is compared with: (a) Q-learning (Qian and Murphy, 2011); (b) Augmented outcome weighted learning (AO-learning) (Liu et al., 2018), which incorporates O-learning with an argumentation term for reducing variability; (c) M-learning (Wu et al., 2020), an alternative policy learning method based on matching; and (d) Policy Tree (Athey and Wager, 2021; Sverdrup et al., 2020), an AIPWE-based policy learning method with fixed-depth decision trees.
[…]
For Policy Tree, we also derive the optimal policy within a class of depth-2 decision trees using R package policytree, where the nuisance functions are estimated by R function double_robust_scores in R package grf.
[…]
For AO-learning, the parametric and non-parametric variants are also considered, represented by the suffixes -Para and -RF respectively. We estimate the propensity score using logistic regression for $W$ on $(1,X)$ using R function glm in the parametric variant, while regression_forest is used for the non-parametric counterpart. The optimal policy is obtained using R package WeightSVM with Gaussian kernel, with the tuning parameters chosen via 5-fold cross-validation. In addition, for the augmentation term defined in Liu et al. (2018), we fit $Y$ on $X$ employing regression_forest.
For M-learning, we establish the matched sets defined in Wu et al. (2020) using Matching and implement the optimization by WeightSVM, with the kernel and tuning parameters chosen following the same configuration as AO-learning. We employ 1-NN matching, following the argument in the simulation studies of Wu et al. (2020). Although the general $M$-NN matching could be considered, our preliminary simulation studies indicate that the simplest 1-NN matching always yields better performance with less computational burden.
For each scenario, we vary the size of training dataset $n\in\{200,500,1000\}$. The criterion to assess the performance is the empirical value function $\mathbb{P}_{n}[Y(\pi(X))]$, where $\mathbb{P}_{n}$ is the empirical average over an independent testing dataset with size $20000$. Each simulation setup is replicated 200 times.
4.3 Simulation Results
Figure 1: Boxplot of empirical value functions in different scenarios, where the main effect model is linear and the contrast function is tree. The global optimal value function is 1.25.
We present the simulation results in this subsection. Figure 1 illustrates the performance in scenarios where the main effect model is linear and the contrast function is tree. The global optimal value function $V(\pi^{*}_{opt})$ is 1.25 in this setup. All the approaches exhibit more satisfactory performance with a larger sample size. As expected, the six variants of MB-learning are competitive since the policy $\pi^{*}_{opt}$ is included in the class $\Pi$. Among them, MB-LR-M5 and MB-LASSO-M5 effectively accommodate the main effect model and thus yield the similarly best performance, demonstrating the necessity of bias correction. The results indicate that MB-learning with a larger number of matches $M$ performs slightly better. However, preliminary simulations suggest that MB-learning is insensitive to $M$. Notably, Policy Tree yields the best performance among the remaining methods. Nevertheless, while Policy Tree behaves comparably as MB-learning in Scenarios 1 and 2, it deteriorates in Scenarios 3-5 due to the extreme propensity scores that introduce significant variability. These phenomena align with previous discussions in the literature regarding approaches based on AIPWE (e.g., Kang and Schafer, 2007; Zhao et al., 2019). In contrast, the variants of MB-learning mitigate this issue by circumventing the use of propensity scores and thus exhibit appealing robustness.
Figure 2: Boxplot of empirical value functions in different scenarios, where the main effect model is linear and the contrast function is non-tree. The global optimal value function is 1.36.
Figure 2 presents the results with linear main effect model and non-tree contrast function, where the global optimal value function is 1.36. These scenarios explore the impact of misspecifying the restricted policy class. In general, MB-learning with bias correction and Q-learning-Para outperform the remaining methods. Among them, Q-learning-Para dominates the others in Scenarios 1 and 2, which can be explained by the reasonable accommodation for the outcome mean model $\mu(x,w)$. However, in Scenarios 3 and 5, the variants of MB-learning with bias correction exhibit similarly superior performance, while Q-learning-Para deteriorates in these situations due to the extreme proportion of two treatment groups, particularly with a small sample size ($n=200$). These results demonstrate the competitive performance of MB-learning, which can be attributed to the desirable approximation of fixed-depth decision tree to the global optimal policy $\pi^{*}_{opt}$ and the robustness of matching. In terms of (A)IPWE-based approaches, Policy Tree and AO-learning result in favorable performance in Scenarios 1 and 2, while behaving unsatisfactorily in the remaining settings due to the extreme weights. It is noteworthy that M-learning yields similarly unfavorable performance as MB-learning without bias correction, which can partly be explained by the matching bias.
Figures S1 and S2 in Supplementary Materials display the results with nonlinear main effect model, where the contrast function is tree and non-tree respectively. Generally, MB-LASSO has the best performance. In contrast, MB-LR behaves similarly to MB-LASSO in Scenarios 1, 2, and 4, while resulting in relatively poor performance in Scenarios 3 and 5. This is not surprising. It seems that the conditional bias is relatively minor in the former scenarios, as illustrated by the results of MB-learning without bias correction, which becomes more significant in the latter situations. Hence, bias correction based on linear regression may be insufficient to mitigate the severe conditional bias in Scenarios 3 and 5. In addition, Q-learning-RF is competitive in these settings, outperforming its parametric counterpart as expected. The remaining approaches exhibit performance similar to those in Figures 1 and 2. Notably, a comparison of the results from Figures 2 and S2 indicates that Q-learning-Para outperforms AO-learning-Para when the main effect model is linear, despite the correct specification of logistic regression for propensity score in AO-learning-Para (Figure 2, Scenario 1). However, AO-learning-Para performs slightly better in general when the main effect is nonlinear, including the situations with incorrectly imposed propensity score model (Figure S2, Scenarios 2 and 5, $n=1000$). One possible explanation for these phenomena is that the performance of these two approaches depends on the extent of model misspecification in outcome mean model and propensity score model, which aligns with the findings discussed in Zhao et al. (2019).
To further illustrate the promising performance of MB-learning, we focus on the comparison with Policy Tree with $n$ varies from 200 to 8000. The results are displayed in Figures S3 and S4 in Supplementary Materials. MB-learning with bias correction outperforms Policy Tree by a large margin when the sample size is small, whereas the performance becomes nearly identical with a larger sample size. This phenomenon can be attributed to the cross-fitting technique employed in Policy Tree, where the fit on nuisance functions by random forest may be inadequate with a small sample size. Similar observations have been mentioned in Wu et al. (2022). Furthermore, this finding provides evidence to support the result of convergence rate discussed in Section 3.
5 Application: Treatment Allocation on NSW Program
The NSW Program aimed to provide work experience for job-seekers facing economic and social difficulties before enrollment. The individuals were randomly assigned to either job training program exposed group or unexposed group. The causal effect of the training program on post-treatment earnings has been analyzed in many literature (e.g., LaLonde, 1986; Dehejia and Wahba, 1999, 2002; Abadie and Imbens, 2011; Imai and Ratkovic, 2014). We utilize the NSW datasets, available at https://users.nber.org/ rdehejia/nswdata2.html, to demonstrate the performance of MB-learning in identifying the optimal policy.
Table 1: The normalized differences of covariates in different datasets.
Covariate
Dataset Age Education Black Hispanic Married No degree RE74 RE75
DW 0.107 0.141 0.044 -0.175 0.094 -0.304 -0.002 0.084
DW-CPS3 -0.128 0.074 0.930 -0.240 -0.426 0.047 -0.378 -0.148
We first implement the proposed method on the dataset used in Dehejia and Wahba (1999) (DW), which is a randomized experiment including 185 training program exposed ($W=1$) and 260 unexposed ($W=0$). The 8-dimensional pre-treatment covariates $X$ are considered, including age, education, black (1 if black, 0 otherwise), Hispanic (1 if Hispanic, 0 otherwise), married (1 if married, 0 otherwise), no degree (1 if no degree, 0 otherwise), earnings in 1974 (RE74), and earnings in 1975 (RE75). The outcome of interest $Y$ is earnings in 1978 (RE78). To assess the covariates balance between the groups, we calculate the normalized difference (Abadie and Imbens, 2011), defined as $(\bar{X}_{1}-\bar{X}_{0})/\sqrt{(S^{2}_{1}+S^{2}_{0})/2}$, where $\bar{X}_{w}=\sum_{i:W_{i}=w}X_{i}/n_{w}$, $S^{2}_{w}=\sum_{i:W_{i}=w}(X_{i}-\bar{X}_{w})^{2}/(n_{w}-1)$, and $n_{w}$ represents the number of units in group $W=w$. The normalized difference values are displayed in Table 1, indicating no significant differences in covariates distributions between the groups. We proceed to derive the optimal policy $\hat{\pi}$ that assigns personalized treatments to the job-seekers for maximizing the post-treatment earnings. Due to legal constraints regarding the use of ethnicity for treatment allocation, black and Hispanic are treated as confounders and are thus excluded from identifying the optimal policy.
We apply the six variants of MB-learning discussed in Section 4 to DW dataset and focus on the comparison with Policy Tree, where the optimal policies are estimated within depth-2 decision trees. The estimated policy $\hat{\pi}$ is evaluated by an AIPWE of the value function $V(\hat{\pi})$:
$\mathbb{P}_{n}\left[\frac{YI\left\{W=\hat{\pi}(X)\right\}}{\tilde{e}(\hat{\pi}(X),X)}-\frac{I\left\{W=\hat{\pi}(X)\right\}-\tilde{e}(\hat{\pi}(X),X)}{\tilde{e}(\hat{\pi}(X),X)}\tilde{\mu}(X,\hat{\pi}(X))\right],$
where $\tilde{e}(w,x)$ and $\tilde{\mu}(x,w)$ represent the estimators of $e(w,x)=\Pr(W=w|X=x)$ and $\mu(x,w)$ respectively. Notably, since this dataset is collected from a randomized experiment, $\tilde{e}(w,x)$ is attained by the empirical proportion of the exposed or unexposed group, while $\tilde{\mu}(x,w)$ is derived by random forest. Following the argument in Zhao et al. (2014), we leverage a cross-validated analysis to assess the performance of the estimated policy. Specifically, the dataset is randomly split into five parts with equal sample sizes, with four parts used for estimating the optimal policy and the remaining fifth part used for evaluating the performance. The training and testing sets are permuted five times such that each part is used as the testing dataset in turn. Cross-validated values for each approach are obtained by averaging the results across the five testing sets. This procedure is repeated 100 times. In addition, to further highlight the improvement of personalized treatments, we also explore the one-size-fits-all policy that all units are either exposed or unexposed, where we calculate the AIPWE-adjusted mean.
Table 2 reports the cross-validated results of the value functions, where higher values are desirable. The results indicate that always being exposed is the most beneficial policy for the job-seekers. One possible explanation is that the job training program generally increases earnings for this population, which implies limited heterogeneity. This phenomenon has also been discussed in Zhou et al. (2023). The results can also be attributed to the model misspecification of the restricted policy class, which may lead to the deterioration of MB-learning and Policy Tree, as discussed in Section 4. Among the personalized policies, MB-LASSO-M5 exhibits the best performance, while the other variants of MB-learning are competitive. In contrast, Policy Tree performs the worst in terms of the cross-validated values, which can partly be explained by the unsatisfactory cross-fitting performance in a relatively small sample. Additionally, the optimal policies estimated by MB-LASSO-M5 and Policy Tree are provided in Figure S5 in Supplementary Materials.
Moreover, we investigate the performance of MB-learning on observational data. The DW dataset is composited with additional 429 unexposed units from Current Population Survey (CPS3), resulting in a synthetic observational study (DW-CPS3). The normalized difference values presented in Table 1 suggest that the discrepancy for covariates between the exposed and unexposed group increases, particularly in race, marital status, and earnings. We conduct cross-validated analysis in this observational study, where both $\tilde{e}(w,x)$ and $\tilde{\mu}(x,w)$ are attained by random forest. As shown in Table 2, the variants of MB-learning have superior performance, with MB-M5 yielding the highest cross-validated value and the smallest standard deviation. It is noteworthy that MB-learning slightly deteriorates after bias correction. One possible reason is the unsatisfactory fit for $\mu(x,w)$, which could be highly nonlinear in this population. In addition, the results illustrate that the individualized policies learned by MB-learning outperform the universal policy of always being exposed. These results are likely to be related to the presence of heterogeneity in this composite dataset, given that CPS3 samples are nationally representative, collected across the United States (Smith and Todd, 2005). The job training program may not universally benefit these composite samples, as they were drawn from different local labor markets. On the other hand, Policy Tree is dominated by the variants of MB-learning, exhibiting the highest variability and yielding a cross-validated value worse than always being exposed. We also present the optimal policies learned by MB-M5 and Policy Tree in Figure S6 in Supplementary Materials.
Table 2: Mean cross-validated value functions in different datasets. The standard deviation is in the parenthesis. In DW dataset, the AIPWEs of the value functions relative to “always being exposed" and “always being unexposed" is 6244.9 and 4617.5, respectively. In DW-CPS3 dataset, the AIPWEs of the value functions relative to “always being exposed" and “always being unexposed" is 6614.2 and 5781.6, respectively.
Policy learning method
Dataset MB-M1 MB-M5 MB-LR-M1 MB-LR-M5 MB-LASSO-M1 MB-LASSO-M5 Policy Tree
DW 5936.7 (220.5) 6036.4 (199.8) 5898.3 (263.8) 5956.0 (198.8) 5916.7 (214.8) 6038.7 (226.2) 5839.8 (226.4)
DW-CPS3 6725.4 (167.5) 6959.0 (127.8) 6757.1 (154.0) 6909.7 (163.4) 6719.0 (159.8) 6949.8 (135.8) 6594.6 (206.0)
6 Discussion
In this paper, we proposed MB-learning for learning the optimal policy in observational studies, leveraging matching estimators in causal inference literature. Specifically, we developed a (bias-corrected) matching-based estimator of the advantage function and estimated the optimal policy by directly maximizing the resulting estimate over a class of policies. The proposed method possesses robustness in observational studies by circumventing the use of inverse probabilities. Theoretically, we established a finite sample bound for the regret of the learned policy, providing rigorous performance guarantee for MB-learning. Empirically, we demonstrated the competitive performance of the proposed method, particularly when dealing with extreme weights or relatively small sample sizes.
There are several interesting directions this work may be extended. First, MB-learning is intended for low-dimensional settings with binary treatment. In causal inference literature, Antonelli et al. (2018) proposed a doubly robust matching estimator for handling high-dimensional covariates, leveraging variable selection techniques for matching with estimated propensity score and prognostic score. In addition, Wu et al. (2022) developed a matching estimator of average causal exposure-response function based on generalized propensity score. These estimators offer inspiration for adapting MB-learning to different scenarios. Second, we developed MB-learning with a fixed number of matches $M$, and our discussion in Section 3 demonstrated that a larger $M$ yields a lower regret bound. Lin et al. (2023) established a remarkable result that bias-corrected matching estimator (Abadie and Imbens, 2011) with a diverging $M$ exhibits doubly robustness and semiparametrically efficiency under regularity conditions. Hence, developing a framework for MB-learning with a diverging $M$ could yield a robust policy learning approach with an improved regret bound. Finally, the robustness of MB-learning can be attributed to the covariates balancing between the groups, which is achieved through directly matching on covariates. It would be of considerable interest to integrate more sophisticated covariate balancing techniques for learning the optimal policy (e.g., Wong and Chan, 2018; Wang and Zubizarreta, 2020; Fan et al., 2022).
Supplementary Materials
Appendix A Technical Proofs
Before the proofs, we introduce some notations here. Let $(\Theta,d)$ denote a semimetric space. For any $\epsilon>0$, let $N(\epsilon,\Theta,d)$ denote the covering number of $\Theta$, i.e., the minimal number of $d$-balls with radius $\epsilon$ needed to cover $\Theta$. We call a set of points is $\epsilon$-separated if the distance between every pair of points is strictly larger than $\epsilon$, and let $D(\epsilon,\Theta,d)$ denote the packing number of $\Theta$, i.e., the maximum number of points in $\Theta$ such that they are $\epsilon$-separated. The entropy numbers are the logarithms of covering and packing numbers respectively. These definitions can be found elsewhere (e.g., van der Vaart and Wellner, 2023). Throughout the proofs, we use $C$ to denote a generic positive constant, which may change from place to place.
A.1 Proof of Lemma 1
Proof.
Let $\mathbf{Y}=(Y_{1},\ldots,Y_{n})^{\top}$, note that
$\hat{Y}(1)-\hat{Y}(0)=(2W_{i}-1)\left(Y_{i}-\frac{1}{M}\underset{j\in\mathcal{J}_{M}(i)}{\sum}Y_{j}\right)=(2W_{i}-1)\left(Y_{i}-\frac{1}{M}M^{\top}_{i}\mathbf{Y}\right),$
and that
$\sum_{i=1}^{n}\left(2\pi(X_{i})-1\right)(2W_{i}-1)M^{\top}_{i}\mathbf{Y}=$ $\sum_{i=1}^{n}\left(2\pi(X_{i})-1\right)(2W_{i}-1)\sum_{l=1}^{n}M_{il}Y_{l}$
$=$ $\sum_{l=1}^{n}\left[\sum_{i=1}^{n}\left(2\pi(X_{i})-1\right)(2W_{i}-1)M_{il}\right]Y_{l}$
$=$ $\sum_{l=1}^{n}\left[\underset{i:M_{il}=1}{\sum}\left(2\pi(X_{i})-1\right)(1-2W_{l})\right]Y_{l}$
$=$ $-\sum_{l=1}^{n}(2W_{l}-1)K_{M}(\pi,l)Y_{l},$
where the third equality follows from the fact that, if $M_{il}=1$, then $l\in\mathcal{J}_{M}(i)$ and thus we have $W_{i}=1-W_{l}$ by the definition of $\mathcal{J}_{M}(i)$. Combining these results, we have
$\hat{A}_{match}(\pi)=$ $\frac{1}{n}\sum_{i=1}^{n}\left(2\pi(X_{i})-1\right)\left(\hat{Y}_{i}(1)-\hat{Y}_{i}(0)\right)$
$=$ $\frac{1}{n}\sum_{i=1}^{n}\left(2\pi(X_{i})-1\right)(2W_{i}-1)\left(Y_{i}-\frac{1}{M}M^{\top}_{i}\mathbf{Y}\right)$
$=$ $\frac{1}{n}\sum_{i=1}^{n}(2W_{i}-1)\left[\left(2\pi(X_{i})-1\right)+\frac{K_{M}(\pi,i)}{M}\right]Y_{i}.$
A.2 Proof of Lemma 2
Proof.
Note that $|K_{M}(\pi,i)|\leq\sum_{j=1}^{n}M_{ji}=K_{M}(i)$, it follows directly from the Lemma 3 in Abadie and Imbens (2006). ∎
A.3 Proof of Lemma 3
Proof.
Using the same argument in Abadie and Imbens (2002), we define two $n\times n$ matrices $\mathbf{G}_{1}$ and $\mathbf{G}_{0}$:
$\mathbf{G}_{1,ij}=\begin{cases}1&\text{ if }W_{i}=1\text{ and }i=j,\\ \frac{1}{M}&\text{ if }W_{i}=0\text{ and }j\in\mathcal{J}_{M}(i),\\ 0&\text{ otherwise},\end{cases}\text{ and }\mathbf{G}_{0,ij}=\begin{cases}1&\text{ if }W_{i}=0\text{ and }i=j,\\ \frac{1}{M}&\text{ if }W_{i}=1\text{ and }j\in\mathcal{J}_{M}(i),\\ 0&\text{ otherwise}.\end{cases}$
Let $\mathbf{\hat{Y}}(1)=\left(\hat{Y}_{1}(1),\ldots,\hat{Y}_{n}(1)\right)^{\top}$ and $\mathbf{\hat{Y}}(0)=\left(\hat{Y}_{1}(0),\ldots,\hat{Y}_{n}(0)\right)^{\top}$ denote the $n\times 1$ vectors of matching estimators for potential outcomes under each treatment, then we have
$\mathbf{\hat{Y}}(1)=\mathbf{G}_{1}\mathbf{Y}\text{ and }\mathbf{\hat{Y}}(0)=\mathbf{G}_{0}\mathbf{Y}.$
Furthermore, let $\bm{\mu}\mathbf{(X\text{,}W)}=\left(\mu(X_{1},W_{1}),\ldots,\mu(X_{n},W_{n})\right)^{\top}$, we have
$\mathbf{G_{1}}\bm{\mu}\mathbf{(X\text{,}W)}=\mathbf{G_{1}}\bm{\mu}\mathbf{(X\text{,}1)}\text{ and }\mathbf{G_{0}}\bm{\mu}\mathbf{(X\text{,}W)}=\mathbf{G_{0}}\bm{\mu}\mathbf{(X\text{,}0)}$ (6)
by the definition of $\mathbf{G}_{1}$ and $\mathbf{G}_{0}$.
Let $\bm{\phi}$ be the $n\times 1$ vector taking value in $\{-1,1\}^{n}$, whose $i$th element is $12\pi(X_{i})-$ 1, and $\bm{\epsilon}$ be the $n\times 1$ with the $i$th element $\epsilon_{i}$ equal to $Y_{i}-\mu(X_{i},W_{i})$. Then we have
$\hat{A}_{match}(\pi)=$ $\frac{1}{n}\bm{\phi}^{\top}(\mathbf{G_{1}}-\mathbf{G_{0}})\mathbf{Y}$ (7)
$=$ $\frac{1}{n}\bm{\phi}^{\top}(\mathbf{G_{1}}-\mathbf{G_{0}})\bm{\mu}\mathbf{(X\text{,}W)}+\frac{1}{n}\bm{\phi}^{\top}(\mathbf{G_{1}}-\mathbf{G_{0}})\bm{\epsilon}$
$=$ $\frac{1}{n}\bm{\phi}^{\top}\mathbf{G_{1}}\bm{\mu}\mathbf{(X\text{,}1)}-\frac{1}{n}\bm{\phi}^{\top}\mathbf{G_{0}}\bm{\mu}\mathbf{(X\text{,}0)}+\frac{1}{n}\bm{\phi}^{\top}(\mathbf{G_{1}}-\mathbf{G_{0}})\bm{\epsilon}$
$=$ $\frac{1}{n}\bm{\phi}^{\top}\left[\bm{\mu}\mathbf{(X\text{,}1)}-\bm{\mu}\mathbf{(X\text{,}0)}\right]+\frac{1}{n}\bm{\phi}^{\top}(\mathbf{G_{1}}-\mathbf{G_{0}})\bm{\epsilon}$
$+\frac{1}{n}\bm{\phi}^{\top}(\mathbf{G_{1}}-\mathbf{I_{n}})\bm{\mu}\mathbf{(X\text{,}1)}-\frac{1}{n}\bm{\phi}^{\top}(\mathbf{G_{0}}-\mathbf{I_{n}})\bm{\mu}\mathbf{(X\text{,}0)},$
where the third equality follows form (6) and $\mathbf{I_{n}}$ is the $n\times n$ identity matrix. Note that the elements of the two $n\times n$ matrices $\mathbf{G_{1}}-\mathbf{I_{n}}$ and $\mathbf{G_{0}}-\mathbf{I_{n}}$ are given by
$(\mathbf{G_{1}}-\mathbf{I_{n}})_{ij}=\begin{cases}-1&\text{ if }W_{i}=0\text{ and }i=j,\\ \frac{1}{M}&\text{ if }W_{i}=0\text{ and }j\in\mathcal{J}_{M}(i),\\ 0&\text{ otherwise},\end{cases}\text{ and }(\mathbf{G_{0}}-\mathbf{I_{n}})_{ij}=\begin{cases}-1&\text{ if }W_{i}=1\text{ and }i=j,\\ \frac{1}{M}&\text{ if }W_{i}=1\text{ and }j\in\mathcal{J}_{M}(i),\\ 0&\text{ otherwise}.\end{cases}$
For each row $i$ such that $W_{i}=0$, we have $(\mathbf{G_{1}}-\mathbf{I_{n}})_{i\cdot}\bm{\mu}\mathbf{(X\text{,}1)}=\frac{1}{M}\sum_{j\in\mathcal{J}_{M}(i)}\mu(X_{j},1)-\mu(X_{i},1)$, where $(\mathbf{G_{1}}-\mathbf{I_{n}})_{i\cdot}$ is the $i$th row of $(\mathbf{G_{1}}-\mathbf{I_{n}})$; for each row $i$ such that $W_{i}=1$, we have $(\mathbf{G_{1}}-\mathbf{I_{n}})_{i\cdot}\bm{\mu}\mathbf{(X\text{,}1)}=0$. Hence, we have
$\bm{\phi}^{\top}(\mathbf{G_{1}}-\mathbf{I_{n}})\bm{\mu}\mathbf{(X\text{,}1)}=\underset{i:W_{i}=0}{\sum}\left(2\pi(X_{i})-1\right)\left[\frac{1}{M}\underset{j\in\mathcal{J}_{M}(i)}{\sum}\mu(X_{j},1)-\mu(X_{i},1)\right],$
and
$\bm{\phi}^{\top}(\mathbf{G_{0}}-\mathbf{I_{n}})\bm{\mu}\mathbf{(X\text{,}0)}=\underset{i:W_{i}=1}{\sum}\left(2\pi(X_{i})-1\right)\left[\frac{1}{M}\underset{j\in\mathcal{J}_{M}(i)}{\sum}\mu(X_{j},0)-\mu(X_{i},0)\right]$
by the similar argument. Furthermore, it follows that
$\bm{\phi}^{\top}(\mathbf{G_{1}}-\mathbf{I_{n}})\bm{\mu}\mathbf{(X\text{,}1)}-$ $\bm{\phi}^{\top}(\mathbf{G_{0}}-\mathbf{I_{n}})\bm{\mu}\mathbf{(X\text{,}0)}$ (8)
$=$ $\sum^{n}_{i=1}(2W_{i}-1)\left(2\pi(X_{i})-1\right)\frac{1}{M}\underset{j\in\mathcal{J}_{M}(i)}{\sum}\left(\mu(X_{i},1-W_{i})-\mu(X_{j},1-W_{i})\right)$
After some algebra and using the similar argument in the proof of Lemma 1, we have
$\bm{\phi}^{\top}\left[\bm{\mu}\mathbf{(X\text{,}1)}-\bm{\mu}\mathbf{(X\text{,}0)}\right]=\sum^{n}_{i=1}\left(2\pi(X_{i})-1\right)\left(\mu(X_{i},1)-\mu(X_{i},0)\right),$ (9)
and
$\bm{\phi}^{\top}(\mathbf{G_{1}}-\mathbf{G_{0}})\bm{\epsilon}=\sum^{n}_{i=1}(2W_{i}-1)\left[(2\pi(X_{i})-1)+\frac{K_{M}(\pi,i)}{M}\right]\epsilon_{i}.$ (10)
Combining (7), (8), (9) and (10), we complete the proof. ∎
A.4 Proof of Lemma 4
Proof.
To prove the lemma, we need to utilize the order of the distances between unit $i$ and its matches. Let $j_{m}(i),m=1,\ldots,M$ denote the $m$th match to unit $i$, then we have $\mathcal{J}_{M}(i)=\left\{j_{1}(i),\ldots,j_{M}(i)\right\}$. It follows that $W_{j_{m}(i)}=1-W_{i}$ and that
$\underset{k:W_{k}=1-W_{i}}{\sum}I\left\{\|X_{k}-X_{i}\|\leq\|X_{j_{m}(i)}-X_{i}\|\right\}=m,$
where, as in Abadie and Imbens (2006), we ignore the ties in matching. We reexpress $B_{M}(\pi)$ in (2) using the indices $\{j_{m}(i),m=1,\ldots,M;i=1,\ldots,n\}$ as follows:
$B_{M}(\pi)=$ $\frac{1}{n}\sum^{n}_{i=1}(2W_{i}-1)\left(2\pi(X_{i})-1\right)\frac{1}{M}\sum_{m=1}^{M}\left[\mu(X_{i},1-W_{i})-\mu(X_{j_{m}(i)},1-W_{i})\right]$ (11)
$=$ $\frac{1}{nM}\sum_{i=1}^{n}\sum_{m=1}^{M}\left(2\pi(X_{i})-1\right)\left\{W_{i}\left[\mu(X_{i},0)-\mu(X_{j_{m}(i)},0)\right]-(1-W_{i})\left[\mu(X_{i},1)-\mu(X_{j_{m}(i)},1)\right]\right\}$
$=$ $\frac{1}{nM}\sum_{i=1}^{n}\sum_{m=1}^{M}B^{i,m}_{M}(\pi),\text{ say}.$
Under the assumption that $\mu(x,0)$ and $\mu(x,1)$ are Lipschitz, we have $\left|B^{i,m}_{M}(\pi)\right|\leq C\|U_{i,m}\|$ for some positive constant $C$, where $U_{i,m}=X_{i}-X_{j_{m}(i)}$ is the matching discrepancy between unit $i$ and $j_{m}(i)$. It follows from Cauchy-Schwarz inequality that
$E\left[n^{2/p}\left(B_{M}(\pi)\right)^{2}\right]\leq n^{2/p}E\left[\frac{1}{nM}\sum_{i=1}^{n}\sum_{m=1}^{M}\left(B^{i,m}_{M}(\pi)\right)^{2}\right]\leq C^{2}n^{2/p}E\left[\frac{1}{n}\sum_{i=1}^{n}\|U_{i,M}\|^{2}\right].$
Using the same argument in Abadie and Imbens (2006, Proof of Theorem 1, Page 258), we have $C^{2}n^{2/p}E\left[\frac{1}{n}\sum_{i=1}^{n}\|U_{i,M}\|^{2}\right]=O(1)$. Therefore, we have $B_{M}(\pi)=O_{p}\left(n^{-1/p}\right)$ by Markov’s inequality.
A.5 Proof of Theorem 1
A.5.1 Proof of Theorem 1 (i)
Proof.
From the result in Lemma 3, we have
$\hat{A}_{match}(\pi)-A(\pi)=\left(\bar{A}(\pi)-A(\pi)\right)+E_{M}(\pi)+B_{M}(\pi).$
We analyze each of the three terms respectively. For the first term, under the Assumption that $\mu(x,w)$ is Lipschitz on a bounded set, then $\sup_{x,w}\mu(x,w)$ is bounded. By week law of large number (van der Vaart, 1998, Proposition 2.16), $\bar{A}(\pi)-A(\pi)\stackrel{{\scriptstyle p}}{{\longrightarrow}}0$. For the second term, let $\bar{\sigma}^{2}=\sup_{x,w}\sigma^{2}(x,w)$, it follows form the Lipschitz assumption that $\bar{\sigma}^{2}$ is bounded. Note that
$E\left[(\sqrt{n}E_{M}(\pi))^{2}\right]=$ $\frac{1}{n}E\left[\sum^{n}_{i=1}(2W_{i}-1)\left[(2\pi(X_{i})-1)+\frac{K_{M}(\pi,i)}{M}\right]\epsilon_{i}\right]^{2}$
$=$ $\frac{1}{n}\sum_{i=1}^{n}E\left[\left(2\pi(X_{i})-1+\frac{K_{M}(\pi,i)}{M}\right)^{2}\epsilon_{i}^{2}\right]$
$=$ $E\left[\left(2\pi(X_{i})-1+\frac{K_{M}(\pi,i)}{M}\right)^{2}\sigma^{2}(X_{i},W_{i})\right]$
$\leq$ $E\left[\left(1+\frac{K_{M}(i)}{M}\right)^{2}\sigma^{2}(X_{i},W_{i})\right]=O(1),$
where the second equality follows from that for $i\neq j$, $E\left[\epsilon_{i}\epsilon_{j}|\{X_{i},W_{i}\}_{i=1}^{n}\right]=0$; the inequality follows from triangle inequality and $|K_{M}(\pi,i)|\leq K_{M}(i)$; and the last equality follows from the boundness of $\bar{\sigma}^{2}$ and Lemma 3 in Abadie and Imbens (2006). Hence, we have $E_{M}(\pi)=O_{p}\left(n^{-1/2}\right)=o_{p}(1)$ by Markov’s inequality. For the third term, it follows from Lemma 4 that $B_{M}(\pi)=O_{p}\left(n^{-1/p}\right)=o_{p}(1)$.
A.5.2 Proof of Theorem 1 (ii)
Proof.
Since
$\sqrt{n}\left(\hat{A}_{match}(\pi)-B_{M}(\pi)-A(\pi)\right)=\sqrt{n}\left(\bar{A}(\pi)-A(\pi)\right)+\sqrt{n}E_{M}(\pi),$
we consider the asymptotic distributions for each term separately. For the first term, note that
$\left(2\pi(X)-1\right)\left(\mu(X,1)-\mu(X,0)\right)-A(\pi)$
has mean zero and finite variance by the assumptions. Therefore, it follows from central limit theorem (van der Vaart, 1998, Proposition 2.17) that
$\sqrt{n}\left(\bar{A}(\pi)-A(\pi)\right)\stackrel{{\scriptstyle d}}{{\rightarrow}}N\left(0,V^{\tau(X)}(\pi)\right).$
For the second term, we consider the similar argument in the proof of Abadie and Imbens (2006, Theorem 4). Let $E_{M,i}(\pi)=(2W_{i}-1)\left[(2\pi(X_{i})-1)+\frac{K_{M}(\pi,i)}{M}\right]\epsilon_{i}$. Conditional on $\{X_{i},W_{i}\}_{i=1}^{n}$, $K_{M}(\pi,i)$ is deterministic. Hence, the unit-level terms $\{E_{M,i}(\pi)\}^{n}_{i=1}$ are independently and not identically distributed with mean zero. We proceed to proof the asymptotic normality of $\sqrt{n}E_{M}(\pi)/\sqrt{V^{E}(\pi)}=\sum_{i=1}^{n}E_{M,i}(\pi)\sqrt{nV^{E}(\pi)}$ using Lindeberg-Feller central limit theorem (Billingsley, 1995, Theorem 27.2). Specifically, given $\{X_{i},W_{i}\}_{i=1}^{n}$, we need to verify Lindeberg condition:
$\frac{1}{nV^{E}(\pi)}\sum_{i=1}^{n}E\left[\left(E_{M,i}(\pi)\right)^{2}I\left\{\left|E_{M,i}(\pi)\right|>\eta\sqrt{nV^{E}(\pi)}\right\}\mid\{X_{i},W_{i}\}_{i=1}^{n}\right]\rightarrow 0,\text{ for all }\eta>0.$
Note that
$E$ $\left[\left(E_{M,i}(\pi)\right)^{2}I\left\{\left|E_{M,i}(\pi)\right|>\eta\sqrt{nV^{E}(\pi)}\right\}|\{X_{i},W_{i}\}_{i=1}^{n}\right]$
$\leq E\left[\left(E_{M,i}(\pi)\right)^{4}\mid\{X_{i},W_{i}\}_{i=1}^{n}\right]^{\frac{1}{2}}E\left[I\left\{\left|E_{M,i}(\pi)\right|>\eta\sqrt{nV^{E}(\pi)}\right\}\mid\{X_{i},W_{i}\}_{i=1}^{n}\right]^{\frac{1}{2}}$
$\leq E\left[\left(E_{M,i}(\pi)\right)^{4}\mid\{X_{i},W_{i}\}_{i=1}^{n}\right]^{\frac{1}{2}}\Pr\left(\left|E_{M,i}(\pi)\right|>\eta\sqrt{nV^{E}(\pi)}\mid\{X_{i},W_{i}\}_{i=1}^{n}\right)$
$\leq E\left[\left(E_{M,i}(\pi)\right)^{4}\mid\{X_{i},W_{i}\}_{i=1}^{n}\right]^{\frac{1}{2}}\frac{E\left[\left(E_{M,i}(\pi)\right)^{2}\mid\{X_{i},W_{i}\}_{i=1}^{n}\right]}{n\eta^{2}V^{E}(\pi)},$
where the first inequality follows from Hölder’s inequality and the third inequality follows from Markov’s inequality.
Let $\bar{c}=\sup_{x,w}E\left[\epsilon^{4}|X=x,W=w\right]<\infty$ by assumption. Since $|K_{M}(\pi,i)|\leq|K_{M}(i)|$, it follows that $|E_{M,i}(\pi)|\leq\left(1+\frac{K_{M}(i)}{M}\right)|\epsilon_{i}|$. In addition, we have $\underline{c}=\inf_{\textbf{X},\textbf{W}}V^{E}(\pi)>0$ by assumption. Therefore, we have
$\frac{1}{nV^{E}(\pi)}$ $\sum_{i=1}^{n}E\left[\left(E_{M,i}(\pi)\right)^{2}I\left\{\left|E_{M,i}(\pi)\right|>\eta\sqrt{nV^{E}(\pi)}\right\}\mid\{X_{i},W_{i}\}_{i=1}^{n}\right]$
$\leq\frac{1}{nV^{E}(\pi)}\sum_{i=1}^{n}\left[\left(1+\frac{K_{M}(i)}{M}\right)^{4}E\left[\epsilon_{i}^{4}\mid\{X_{i},W_{i}\}_{i=1}^{n}\right]\right]^{\frac{1}{2}}\frac{\left(1+\frac{K_{M}(i)}{M}\right)^{2}\sigma^{2}(X_{i},W_{i})}{n\eta^{2}V^{E}(\pi)}$
$\leq\frac{\bar{c}^{1/2}\bar{\sigma}^{2}}{\eta^{2}{\underline{c}}^{2}}\frac{1}{n}\left[\frac{1}{n}\sum_{i=1}^{n}\left(1+\frac{K_{M}(i)}{M}\right)^{4}\right].$
It follows from the Lemma 3 of Abadie and Imbens (2006) that $E\left[\left(1+\frac{K_{M}(i)}{M}\right)^{4}\right]$ is bounded. Therefore, we have
$\frac{1}{nV^{E}(\pi)}\sum_{i=1}^{n}E\left[\left(E_{M,i}(\pi)\right)^{2}I\left\{\left|E_{M,i}(\pi)\right|>\eta\sqrt{nV^{E}(\pi)}\right\}\mid\{X_{i},W_{i}\}_{i=1}^{n}\right]\stackrel{{\scriptstyle p}}{{\rightarrow}}0,$
and thus $\sqrt{n}E_{M}(\pi)/\sqrt{V^{E}(\pi)}\stackrel{{\scriptstyle d}}{{\rightarrow}}N(0,1)$.
Note that $\sqrt{n}\left(\bar{A}(\pi)-A(\pi)\right)$ and $\sqrt{n}E_{M}(\pi)/\sqrt{V^{E}(\pi)}$ are asymptotically independent. Using the same argument in Abadie and Imbens (2006, Proof of Theorem 4, Page 266), we complete the proof. ∎
A.6 Proof of Corollary 2
Proof.
We only need to verify that $\sqrt{n}\left(B_{M}(\pi)-\hat{B}_{M}(\pi)\right)=o_{p}(1)$. To do this, we need to use the result of unit-level bias correction (Abadie and Imbens, 2011, Lemma A.2), which is given as follows:
Lemma A.2. Abadie and Imbens (2011). Assume the conditions in Theorem 2 of Abadie and Imbens (2011) hold, then
$\underset{i\leq n}{\max}\left|\mu(X_{i},w)-\mu(X_{j_{m}(i)},w)-\left(\hat{\mu}(X_{i},w)-\hat{\mu}(X_{j_{m}(i)},w)\right)\right|=o_{p}\left(n^{-1/2}\right)$
for $w=0,1$.
It follows from (11) that
$|B_{M}$ $(\pi)-\hat{B}_{M}(\pi)|$
$=$ $\frac{1}{nM}\bigg{|}\sum_{i=1}^{n}\sum_{m=1}^{M}\left(2\pi(X_{i})-1\right)\bigg{\{}W_{i}\bigg{[}\mu(X_{i},0)-\mu(X_{j_{m}(i)},0)-\left(\hat{\mu}(X_{i},0)-\hat{\mu}(X_{j_{m}(i)},0)\right)\bigg{]}$
$-(1-W_{i})\bigg{[}\mu(X_{i},1)-\mu(X_{j_{m}(i)},1)-\left(\hat{\mu}(X_{i},1)-\hat{\mu}(X_{j_{m}(i)},1)\right)\bigg{]}\bigg{\}}\bigg{|}$
$\leq$ $\frac{1}{nM}\sum_{i=1}^{n}\sum_{m=1}^{M}\bigg{[}\left|\mu(X_{i},0)-\mu(X_{j_{m}(i)},0)-\left(\hat{\mu}(X_{i},0)-\hat{\mu}(X_{j_{m}(i)},0)\right)\right|$
$+\left|\mu(X_{i},1)-\mu(X_{j_{m}(i)},1)-\left(\hat{\mu}(X_{i},1)-\hat{\mu}(X_{j_{m}(i)},1)\right)\right|\bigg{]}$
$\leq$ $\frac{1}{M}\sum_{m=1}^{M}\underset{i\leq n}{\max}\underset{w=0,1}{\sum}\left|\mu(X_{i},w)-\mu(X_{j_{m}(i)},w)-\left(\hat{\mu}(X_{i},w)-\hat{\mu}(X_{j_{m}(i)},w)\right)\right|.$
Using Lemma A.2 in Abadie and Imbens (2011), we have $B_{M}(\pi)-\hat{B}_{M}(\pi)=o_{p}\left(n^{-1/2}\right)$.
A.7 Proof of Theorem 3
Proof.
Since $R(\hat{\pi})=\left(A(\pi^{*})-A(\hat{\pi})\right)/2$ by definition, it suffices to analyze $A(\pi^{*})-A(\hat{\pi})$ throughout the proof. Let $\Delta(\pi_{1},\pi_{2})=A(\pi_{1})-A(\pi_{2})$, $\bar{\Delta}(\pi_{1},\pi_{2})=\bar{A}(\pi_{1})-\bar{A}(\pi_{2})$, and $\hat{\Delta}(\pi_{1},\pi_{2})=\hat{A}^{bc}_{match}(\pi_{1})-\hat{A}^{bc}_{match}(\pi_{2})$. It follows from triangle inequality that
$A(\pi^{*})-A(\hat{\pi})=$ $\hat{A}^{bc}_{match}(\pi^{*})-\hat{A}^{bc}_{match}(\hat{\pi})+A(\pi^{*})-A(\hat{\pi})-\left(\hat{A}^{bc}_{match}(\pi^{*})-\hat{A}^{bc}_{match}(\hat{\pi})\right)$
$\leq$ $A(\pi^{*})-A(\hat{\pi})-\left(\hat{A}^{bc}_{match}(\pi^{*})-\hat{A}^{bc}_{match}(\hat{\pi})\right)$
$=$ $\Delta(\pi^{*},\hat{\pi})-\hat{\Delta}(\pi^{*},\hat{\pi})$
$\leq$ $\underset{\pi_{1},\pi_{2}\in\Pi}{\sup}|\Delta(\pi_{1},\pi_{2})-\hat{\Delta}(\pi_{1},\pi_{2})|$
$\leq$ $\underset{\pi_{1},\pi_{2}\in\Pi}{\sup}|\bar{\Delta}(\pi_{1},\pi_{2})-\Delta(\pi_{1},\pi_{2})|+\underset{\pi_{1},\pi_{2}\in\Pi}{\sup}|\bar{\Delta}(\pi_{1},\pi_{2})-\hat{\Delta}(\pi_{1},\pi_{2})|$
$\leq$ $\underset{\pi_{1},\pi_{2}\in\Pi}{\sup}|\bar{\Delta}(\pi_{1},\pi_{2})-\Delta(\pi_{1},\pi_{2})|+2\underset{\pi\in\Pi}{\sup}|B_{M}(\pi)-\hat{B}_{M}(\pi)|+2\underset{\pi\in\Pi}{\sup}|E_{M}(\pi)|,$ (12)
where the last inequality follows from
$\hat{\Delta}(\pi_{1},\pi_{2})-\bar{\Delta}(\pi_{1},\pi_{2})=B_{M}(\pi_{1})-\hat{B}_{M}(\pi_{1})+E_{M}(\pi_{1})-[B_{M}(\pi_{2})-\hat{B}_{M}(\pi_{2})+E_{M}(\pi_{2})].$
We will derive the high probability bounds for the three terms in (A.7) in Step 1-3 separately, and then combine these bounds to complete the proof in Step 4.
A.7.1 Step 1: High probability bound for the first term in (A.7)
To derive this bound, we utilize Lemma 2 in Zhou et al. (2023). First, we rewrite $A(\pi)$ and $\bar{A}(\pi)$ in a vector-based form in order to leverage their Lemma 2. Let
$\Gamma^{vec}_{i}=\left[\begin{aligned} \begin{array}[]{c}\mu(X_{i},0)-\mu(X_{i},1)\\ \mu(X_{i},1)-\mu(X_{i},0)\end{array}\end{aligned}\right]\text{ and }\pi^{vec}(X_{i})=\left[\begin{aligned} \begin{array}[]{c}1-\pi(X_{i})\\ \pi(X_{i})\end{array}\end{aligned}\right],\text{ for }i=1,\ldots,n.$
[…]
$\underset{\pi_{1},\pi_{2}\in\Pi}{\sup}|\bar{\Delta}(\pi_{1},\pi_{2})-\Delta(\pi_{1},\pi_{2})|\leq\left(c_{1}\sqrt{\text{VC}(\Pi)}+c_{2}+\sqrt{2\log\frac{1}{\delta}}\right)\sqrt{\frac{V^{*}}{n}}+o(\frac{1}{\sqrt{n}}),$ (13)
where $V^{*}=4\sup_{\pi_{1},\pi_{2}\in\Pi}E\left[\left(\mu(X,1)-\mu(X,0)\right)^{2}\left(\pi_{1}(X)-\pi_{2}(X)\right)^{2}\right].$
For the last term in (13), let $c_{3}>0$, then there exists $N(c_{3})$ such that for all $n>N(c_{3})$, the last term is upper bounded by $\frac{c_{3}}{\sqrt{n}}$. Therefore, for all $n>N(c_{3})$, with probability at least $1-2\delta$, we have
$\underset{\pi_{1},\pi_{2}\in\Pi}{\sup}|\bar{\Delta}(\pi_{1},\pi_{2})-\Delta(\pi_{1},\pi_{2})|\leq\left(c_{1}\sqrt{\text{VC}(\Pi)}+c_{2}+\sqrt{2\log\frac{1}{\delta}}\right)\sqrt{\frac{V^{*}}{n}}+\frac{c_{3}}{\sqrt{n}}.$ (14)
A.7.2 Step 2: High probability bound for the second term in (A.7)
It follows from the proof of Corollary 2 that
$\sup_{\pi\in\Pi}|B_{M}(\pi)-\hat{B}_{M}(\pi)|\leq\frac{1}{M}\sum_{m=1}^{M}\underset{i\leq n}{\max}\underset{w=0,1}{\sum}\left|\mu(X_{i},w)-\mu(X_{j_{m}(i)},w)-\left(\hat{\mu}(X_{i},w)-\hat{\mu}(X_{j_{m}(i)},w)\right)\right|.$
Using Lemma A.2 in Abadie and Imbens (2011), we have $\sup_{\pi\in\Pi}|B_{M}(\pi)-\hat{B}_{M}(\pi)|=o_{p}\left(\frac{1}{\sqrt{n}}\right)$. Let $c_{4}>0$, for any $\delta>0$, then there exists $N(c_{4},\delta)$ such that for all $n>N(c_{4},\delta)$, with probability at least $1-\delta$,
$2\underset{\pi\in\Pi}{\sup}|B_{M}(\pi)-\hat{B}_{M}(\pi)|\leq\frac{c_{4}}{\sqrt{n}}.$ (15)
A.7.3 Step 3: High probability bound for the third term in (A.7)
By the definition of $E_{M}(\pi)$ and triangle inequality, we have
$\underset{\pi\in\Pi}{\sup}|E_{M}(\pi)|\leq$ $\underset{\pi\in\Pi}{\sup}\left|\frac{1}{n}\sum^{n}_{i=1}(2W_{i}-1)(2\pi(X_{i})-1)\epsilon_{i}\right|+\underset{\pi\in\Pi}{\sup}\left|\frac{1}{n}\sum^{n}_{i=1}(2W_{i}-1)\frac{K_{M}(\pi,i)}{M}\epsilon_{i}\right|$
$\leq$ $\left|\frac{1}{n}\sum^{n}_{i=1}(2W_{i}-1)\epsilon_{i}\right|+\left|\frac{1}{n}\sum^{n}_{i=1}(2W_{i}-1)\frac{K_{M}(i)}{M}\epsilon_{i}\right|+2\underset{\pi\in\Pi}{\sup}\left|\frac{1}{n}\sum^{n}_{i=1}(2W_{i}-1)\pi(X_{i})\epsilon_{i}\right|$
$+\frac{2}{M}\underset{\pi\in\Pi}{\sup}\left|\frac{1}{n}\sum^{n}_{i=1}(2W_{i}-1)\left[\sum_{j=1}^{n}\pi(X_{j})M_{ji}\right]\epsilon_{i}\right|$
[…]
where the second inequality follows from the fact that
$K_{M}(\pi,i)=\sum^{n}_{j=1}\left(2\pi(X_{j})-1\right)M_{ji}=2\sum_{j=1}^{n}\pi(X_{j})M_{ji}-K_{M}(i).$
We will derive high probability upper bounds for these four terms separately.
Step 3.1: Bound the (I) term in (A.7.3)
We bound (I) using Hoeffding’s inequality (Vershynin, 2018, Theorem 2.2.6). Under the assumption that $|Y_{i}|\leq L$, we have $|\epsilon_{i}|=|Y_{i}-E[Y_{i}|X_{i},W_{i}]|\leq 2L$ by triangle inequality. In addition, note that $E[(2W_{i}-1)\epsilon_{i}]=E\left[(2W_{i}-1)E\left[\epsilon_{i}\mid X_{i},W_{i}\right]\right]=0$. By Hoeffding’s inequality, for any $t>0$,
$\Pr\left(\left|\frac{1}{n}\sum_{i=1}^{n}(2W_{i}-1)\epsilon_{i}\right|>\frac{t}{\sqrt{n}}\right)\leq 2\exp\left(\frac{-t^{2}}{8L^{2}}\right).$
Let $\delta=\exp\left(\frac{-t^{2}}{8L^{2}}\right)$, we have
$\Pr\left(\left|\frac{1}{n}\sum_{i=1}^{n}(2W_{i}-1)\epsilon_{i}\right|\leq 2\sqrt{2}L\sqrt{\frac{\log\frac{1}{\delta}}{n}}\right)\geq 1-2\delta.$ (17)
Step 3.2: Bound the (II) term in (A.7.3)
For (II), we also apply Hoeffding’s inequality to derive the high probability bound. Given $\{X_{i},W_{i}\}^{n}_{i=1}$, $K_{M}(i)$ is constant and $(2W_{i}-1)K_{M}(i)\epsilon_{i}$ is independent with mean zero. For any $t>0$, we have
$\Pr\left(\left|\frac{1}{n}\sum_{i=1}^{n}(2W_{i}-1)\frac{K_{M}(i)}{M}\epsilon_{i}\right|>\frac{t}{n}\sqrt{\sum_{i=1}^{n}K^{2}_{M}(i)}\mid\{X_{i},W_{i}\}^{n}_{i=1}\right)\leq 2\exp\left(-\frac{M^{2}t^{2}}{8L^{2}}\right).$
Taking expectation with respect to (w.r.t.) $\{X_{i},W_{i}\}^{n}_{i=1}$, we have
$\Pr\left(\left|\frac{1}{n}\sum_{i=1}^{n}(2W_{i}-1)\frac{K_{M}(i)}{M}\epsilon_{i}\right|>\frac{t}{n}\sqrt{\sum_{i=1}^{n}K^{2}_{M}(i)}\right)\leq 2\exp\left(-\frac{M^{2}t^{2}}{8L^{2}}\right).$
Let $\delta=\exp\left(\frac{-M^{2}t^{2}}{8L^{2}}\right)$, it follows that
$\Pr\left(\left|\frac{1}{n}\sum_{i=1}^{n}(2W_{i}-1)\frac{K_{M}(i)}{M}\epsilon_{i}\right|\leq\frac{2\sqrt{2}L}{Mn}\sqrt{\log\frac{1}{\delta}\sum_{i=1}^{n}K^{2}_{M}(i)}\right)\geq 1-2\delta.$ (18)
Step 3.3: Bound the (III) term in (A.7.3)
To derive this bound, we first analyze the order of $E\left[\sup_{\pi\in\Pi}\left|\frac{1}{n}\sum^{n}_{i=1}(2W_{i}-1)\pi(X_{i})\epsilon_{i}\right|\right]$ using empirical process theory. We proceed to consider the empirical process indexed by the functional class $\Pi$. To simplify our analysis, we assume the cardinality of the class $\Pi$ is finite hereafter. Our results can be extended to a countable $\Pi$ by Monotone Convergence Theorem (Massart, 2007).
Denote the functional class
=$\mathcal{H}=\left\{f_{\pi}(x,w,\epsilon)=(2w-1)\epsilon\pi(x),\pi\in\Pi\right\},$
and the envelope function of $\mathcal{H}$ is $H(x,w,\epsilon)=|\epsilon|$. Therefore, we have $L_{2}(\mathbb{P})$ norm $\|H\|_{P,2}=\sqrt{E\epsilon^{2}}\leq 2L$. Recall that $E[(2W_{i}-1)\pi(X_{i})\epsilon_{i}]=0$, we have
$\underset{\pi\in\Pi}{\sup}\left|\frac{1}{n}\sum^{n}_{i=1}(2W_{i}-1)\pi(X_{i})\epsilon_{i}\right|=\underset{f\in\mathcal{H}}{\sup}|\mathbb{P}_{n}f|=\underset{f\in\mathcal{H}}{\sup}|(\mathbb{P}_{n}-\mathbb{P})f|,$
where $\mathbb{P}_{n}$ and $\mathbb{P}$ denote the empirical measure and the expectation respectively. Note that the functional class $\mathcal{H}$ can be rewritten as the pointwise product between the class $\Pi$ and the function $g(w,\epsilon)=(2w-1)\epsilon$, i.e., $\mathcal{H}=\Pi\cdot g=\{\pi g,\pi\in\Pi\}$. Hence, by the results of VC-class (Kosorok, 2008, Lemma 9.9), $\mathcal{H}$ is VC-class and $\text{VC}(\mathcal{H})\leq 2\text{VC}(\Pi)+1$ (note that Lemma 9.9 in Kosorok (2008) is stated in terms of VC-index, which is equal to the corresponding VC dimension plus one).
Since the norm of the empirical process indexed by a VC-class can be bounded by its VC dimension and the $L_{2}(\mathbb{P})$-seminorm of its envelope function (van der Vaart and Wellner, 2023, Example 3.5.13), then there exists a universal constant $c_{5}>0$ such that
$E\left[\underset{f\in\mathcal{H}}{\sup}|(\mathbb{P}_{n}-\mathbb{P})f|\right]\leq c_{5}L\sqrt{\frac{\text{VC}(\Pi)}{n}}.$ (19)
Second, by Functional Hoeffding Theorem (Wainwright, 2019, Theorem 3.26), for any $t>0$ we have
$\Pr\left(\underset{\pi\in\Pi}{\sup}\left|\frac{1}{n}\sum^{n}_{i=1}(2W_{i}-1)\pi(X_{i})\epsilon_{i}\right|>E\left[\underset{\pi\in\Pi}{\sup}\left|\frac{1}{n}\sum^{n}_{i=1}(2W_{i}-1)\pi(X_{i})\epsilon_{i}\right|\right]+\frac{t}{\sqrt{n}}\right)$ (20)
$\leq\exp\left(-\frac{t^{2}}{64L^{2}}\right).$
Let $\delta=\exp\left(-\frac{t^{2}}{64L^{2}}\right)$, combining (19) and (20) we have
$\Pr\left(\underset{\pi\in\Pi}{\sup}\left|\frac{1}{n}\sum^{n}_{i=1}(2W_{i}-1)\pi(X_{i})\epsilon_{i}\right|\leq c_{5}L\sqrt{\frac{\text{VC}(\Pi)}{n}}+8L\sqrt{\frac{\log\frac{1}{\delta}}{n}}\right)\geq 1-\delta.$ (21)
Step 3.4: Bound the (IV) term in (A.7.3)
To do this, we first derive the order of
$E\left[\underset{\pi\in\Pi}{\sup}\left|\frac{1}{n}\sum^{n}_{i=1}(2W_{i}-1)\left[\sum_{j=1}^{n}\pi(X_{j})M_{ji}\right]\epsilon_{i}\right|\right]$ (22)
using Dudley’s entropy integral bound (Vershynin, 2018, Theorem 8.1.3). Then we consider the concentration inequality of
$\underset{\pi\in\Pi}{\sup}\left|\frac{1}{n}\sum^{n}_{i=1}(2W_{i}-1)\left[\sum_{j=1}^{n}\pi(X_{j})M_{ji}\right]\epsilon_{i}\right|$
around the expectation (22) using Talagrand’s inequality (Massart, 2007). In our proof, we need to use the maximal inequality of the absolute values of sub-Gaussian random variables, which is given as follows:
Corollary 7.2. Zhang and Chen (2021). Let $\{X_{i}\}_{i=1}^{n}$ are sub-Gaussian random variables such that for $t\geq 0$, $\Pr(|X_{i}|>t)\leq 2\exp\left(-t^{2}/(2\sigma^{2})\right)$, then
$E\left[\underset{i\leq n}{\max}|X_{i}|\right]\leq\sigma\sqrt{2\log(2n)}.$ (23)
Step 3.4.1: The order of expectation
We define the unit-level term $a_{i}(\pi)$ in (IV) as
$a_{i}(\pi)=(2W_{i}-1)\left[\sum_{j=1}^{n}\pi(X_{j})M_{ji}\right]\epsilon_{i},\text{ for }i=1,\ldots,n.$
Given $\{X_{i},W_{i}\}^{n}_{i=1}$, $a_{i}(\pi)$ is independently and not identically distributed with $E\left[a_{i}(\pi)\right]=0$. We define a Rademacher process $G^{0}_{n}(\pi)=\frac{1}{\sqrt{n}}\sum^{n}_{i=1}a_{i}(\pi)Z_{i}$, where $\{Z_{i}\}^{n}_{i=1}$ is a sequence of independent Rademacher variables independent of $\{X_{i},W_{i},\epsilon_{i}\}^{n}_{i=1}$. Therefore, conditional on $\{X_{i},W_{i}\}^{n}_{i=1}$, it follows from the symmetrization inequality (Boucheron et al., 2013, Lemma 11.4) that
$E\left[\underset{\pi\in\Pi}{\sup}\left|\frac{1}{n}\sum^{n}_{i=1}(2W_{i}-1)\left[\sum_{j=1}^{n}\pi(X_{j})M_{ji}\right]\epsilon_{i}\right|\right]=$ $\frac{1}{\sqrt{n}}E\left[\underset{\pi\in\Pi}{\sup}\left|\frac{1}{\sqrt{n}}\sum^{n}_{i=1}a_{i}(\pi)\right|\right]$ (24)
$\leq$ $\frac{2}{\sqrt{n}}E\left[\underset{\pi\in\Pi}{\sup}\left|\frac{1}{\sqrt{n}}\sum^{n}_{i=1}a_{i}(\pi)Z_{i}\right|\right]$
$=$ $\frac{2}{\sqrt{n}}E\left[\underset{\pi\in\Pi}{\sup}\left|G^{0}_{n}(\pi)\right|\right].$
We proceed to derive an upper bound for the conditional expectation $E\left[\sup_{\pi\in\Pi}\left|G^{0}_{n}(\pi)\right|\right]$ with fixed $\{X_{i},W_{i}\}^{n}_{i=1}$.
For any function $f:\mathcal{X}\rightarrow\mathbb{R}$, the $L_{2}(\mathbb{P}_{n})$-seminorm is defined as $\|f\|_{n}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}f^{2}(X_{i})}$ (van der Vaart and Wellner, 2023). Given $\{X_{i},W_{i},\epsilon_{i}\}^{n}_{i=1}$, the sequence $\{a_{i}(\pi)\}_{i=1}^{n}$ is deterministic. Therefore, by Hoeffding’s inequality, we have
$\Pr\left(\left|G^{0}_{n}(\pi_{1})-G^{0}_{n}(\pi_{2})\right|\geq t\right)\leq 2\exp\left(-t^{2}\bigg{/}\left\{\frac{2}{n}\sum_{i=1}^{n}\epsilon^{2}_{i}\left[\sum_{j=1}^{n}\left(\pi_{1}(X_{j})-\pi_{2}(X_{j})\right)M_{ji}\right]^{2}\right\}\right),$
for any $\pi_{1},\pi_{2}\in\Pi$ and any $t\geq 0$. Note that
$\frac{1}{n}$ $\sum_{i=1}^{n}\epsilon^{2}_{i}\left[\sum_{j=1}^{n}\left(\pi_{1}(X_{j})-\pi_{2}(X_{j})\right)M_{ji}\right]^{2}=\frac{1}{n}\sum_{i=1}^{n}\epsilon^{2}_{i}\left[\underset{j:M_{ji}=1}{\sum}\left(\pi_{1}(X_{j})-\pi_{2}(X_{j})\right)\right]^{2}$
$\leq$ $\frac{1}{n}\sum_{i=1}^{n}\epsilon^{2}_{i}K_{M}(i)\left[\underset{j:M_{ji}=1}{\sum}\left(\pi_{1}(X_{j})-\pi_{2}(X_{j})\right)^{2}\right]$
$\leq$ $\frac{4L^{2}\underset{i\leq n}{\max}K_{M}(i)}{n}\sum_{i=1}^{n}\left[\underset{j:M_{ji}=1}{\sum}\left(\pi_{1}(X_{j})-\pi_{2}(X_{j})\right)^{2}\right]$
$=$ $\frac{4L^{2}\underset{i\leq n}{\max}K_{M}(i)}{n}\sum_{i=1}^{n}\left[\sum_{j=1}^{n}\left(\pi_{1}(X_{j})-\pi_{2}(X_{j})\right)^{2}M_{ji}\right]$
$=$ $\frac{4L^{2}\underset{i\leq n}{\max}K_{M}(i)}{n}\sum_{j=1}^{n}\left[\left(\pi_{1}(X_{j})-\pi_{2}(X_{j})\right)^{2}\sum_{i=1}^{n}M_{ji}\right]$
$=$ $\frac{4ML^{2}\underset{i\leq n}{\max}K_{M}(i)}{n}\sum_{j=1}^{n}\left[\left(\pi_{1}(X_{j})-\pi_{2}(X_{j})\right)^{2}\right]=4ML^{2}\underset{i\leq n}{\max}K_{M}(i)\|\pi_{1}-\pi_{2}\|^{2}_{n},$
where the first inequality follows from Cauchy-Swarchz inequality and the fact that $K_{M}(i)=\sum_{j=1}^{n}M_{ji}=\sum_{j:M_{ji}=1}1$; the fourth equality follows from $\sum_{i=1}^{n}M_{ji}=M$, i.e., the cardinality of the set $\mathcal{J}_{M}(j)$ is $M$. Therefore, conditional on $\{X_{i},W_{i},\epsilon_{i}\}^{n}_{i=1}$, $G^{0}_{n}(\pi)$ is a sub-Gaussian process with mean zero relative to $L_{2}(\mathbb{P}_{n})$-seminorm $\|\cdot\|_{n}$ in the sense that
$\Pr\left(\left|G^{0}_{n}(\pi_{1})-G^{0}_{n}(\pi_{2})\right|\geq t\right)\leq 2\exp\left(\frac{-t^{2}}{8ML^{2}\underset{i\leq n}{\max}K_{M}(i)\|\pi_{1}-\pi_{2}\|^{2}_{n}}\right),$ (25)
for any $\pi_{1},\pi_{2}\in\Pi$ and any $t\geq 0$.
We will derive an upper bound for $E_{Z}\left[\sup_{\pi\in\Pi}\left|G^{0}_{n}(\pi)\right|\right]$ by slightly modifying the chaining technique in the proof of Dudley’s entropy integral bound, where $E_{Z}(\cdot)$ denotes the expectation w.r.t. $Z$. We follow the similar argument in Sen (2021), though the proof can be found elsewhere (e.g., Vershynin, 2018; Wainwright, 2019; van der Vaart and Wellner, 2023).
Let $\tilde{\pi}_{0}(\cdot)\equiv 0$, we consider the class $\tilde{\Pi}=\Pi\cup\{\tilde{\pi}_{0}\}$. The diameter of $\tilde{\Pi}$ is $\tilde{D}=\sup_{\pi^{\prime},\pi^{\prime\prime}\in\tilde{\Pi}}\|\pi^{\prime}-\pi^{\prime\prime}\|_{n}$, and note that $\tilde{D}\leq 1$. For $k\geq 1$, let $\tilde{\Pi}_{k}$ denote the maximal $\tilde{D}2^{-k}$-separated subset of $\tilde{\Pi}$ and hence the cardinality of $\tilde{\Pi}_{k}$ is $D(2^{-k},\tilde{\Pi},L_{2}(\mathbb{P}_{n}))$. By the maximality, it follows that
$\sup_{\pi^{\prime}\in\tilde{\Pi}}\inf_{\pi^{\prime\prime}\in\tilde{\Pi}_{k}}\|\pi^{\prime}-\pi^{\prime\prime}\|_{n}\leq\tilde{D}2^{-k}.$ (26)
In addition, since $\tilde{\Pi}$ is finite under our assumption, there exists a sufficiently large $S$ such that $\tilde{\Pi}_{S}=\tilde{\Pi}$. For any $\pi\in\tilde{\Pi}$, we can choose $\tilde{\pi}_{k}$ such that $\tilde{\pi}_{k}=\arg\inf_{\psi\in\tilde{\Pi}_{k}}\|\pi-\psi\|_{n}$, and therefore $\|\pi-\tilde{\pi}_{k}\|_{n}\leq\tilde{D}2^{-k}$ by (26). Let $\tilde{\Pi}_{0}=\{\tilde{\pi}_{0}\}$, we also have $\|\pi-\tilde{\pi}_{0}\|_{n}\leq\tilde{D}$. Furthermore, we note that $\tilde{\pi}_{S}=\pi$ since $\tilde{\Pi}_{S}=\tilde{\Pi}$. Finally, for any $\pi\in\tilde{\Pi}$, we can find a sequence of policies $\tilde{\pi}_{0},\tilde{\pi}_{1},\ldots,\tilde{\pi}_{S}$ that approximates $\pi$.
Note that $G^{0}_{n}(\tilde{\pi}_{0})=0$ and that $G^{0}_{n}(\tilde{\pi}_{S})=G^{0}_{n}(\pi)$, we rewrite the process $G^{0}_{n}(\pi)$ as
$G^{0}_{n}(\pi)=\sum_{k=1}^{S}\left(G^{0}_{n}(\tilde{\pi}_{k})-G^{0}_{n}(\tilde{\pi}_{k-1})\right).$
It follows from triangle inequality that
$E_{Z}\left[\underset{\pi\in\Pi}{\sup}\left|G^{0}_{n}(\pi)\right|\right]=E_{Z}\left[\underset{\pi\in\tilde{\Pi}}{\sup}\left|G^{0}_{n}(\pi)\right|\right]\leq\sum_{k=1}^{S}E_{Z}\left[\underset{\pi\in\tilde{\Pi}}{\sup}\left|G^{0}_{n}(\tilde{\pi}_{k})-G^{0}_{n}(\tilde{\pi}_{k-1})\right|\right].$ (27)
Recall that in (25), we have showed that $G^{0}_{n}(\pi)$ is a sub-Gaussian process given $\{X_{i},W_{i},\epsilon_{i}\}^{n}_{i=1}$. Note that $\|\tilde{\pi}_{k}-\tilde{\pi}_{k-1}\|_{n}\leq\|\tilde{\pi}_{k}-\pi\|_{n}+\|\tilde{\pi}_{k-1}-\pi\|_{n}\leq 3\tilde{D}2^{-k}$, we have
$\Pr\left(\left|G^{0}_{n}(\tilde{\pi}_{k})-G^{0}_{n}(\tilde{\pi}_{k-1})\right|\geq t\right)\leq 2\exp\left(\frac{-t^{2}}{72\cdot 2^{-2k}M\tilde{D}^{2}L^{2}\underset{i\leq n}{\max}K_{M}(i)}\right).$
Since the set of random variables $\left\{G^{0}_{n}(\pi^{\prime})-G^{0}_{n}(\pi^{\prime\prime})\right\}_{\pi^{\prime}\in\tilde{\Pi}_{k},\pi^{\prime\prime}\in\tilde{\Pi}_{k-1}}$ contains at most
$D(2^{-(k-1)},\tilde{\Pi},L_{2}(\mathbb{P}_{n}))\cdot D(2^{-k},\tilde{\Pi},L_{2}(\mathbb{P}_{n}))\leq D^{2}(2^{-k},\tilde{\Pi},L_{2}(\mathbb{P}_{n}))$
elements. By the maximal inequality of sub-Gaussian random variables (23), we have
$E_{Z}\left[\underset{\pi\in\tilde{\Pi}}{\sup}\left|G^{0}_{n}(\tilde{\pi}_{k})-G^{0}_{n}(\tilde{\pi}_{k-1})\right|\right]\leq 2^{-k}C\tilde{D}L\sqrt{M\underset{i\leq n}{\max}K_{M}(i)}\sqrt{\log\left(2D(2^{-k},\tilde{\Pi},L_{2}(\mathbb{P}_{n}))\right)},$
and therefore by (27) we have (recall that the generic constant $C$ may varies from place to place)
$E_{Z}\left[\underset{\pi\in\tilde{\Pi}}{\sup}\left|G^{0}_{n}(\pi)\right|\right]\leq$ $\sum_{k=1}^{S}2^{-k}C\tilde{D}L\sqrt{M\underset{i\leq n}{\max}K_{M}(i)}\sqrt{\log\left(2D(2^{-k},\tilde{\Pi},L_{2}(\mathbb{P}_{n}))\right)}$
[…]
where the fourth inequality follows from the fact that, for $\eta\leq\tilde{D}/2$, the packing number $D(\eta,\tilde{\Pi},L_{2}(\mathbb{P}_{n}))\geq 2$; the fifth inequality follows from the relationship between covering number and packing number (van der Vaart and Wellner, 2023, Page 147); and the last inequality follows from the change of variable and the fact that $\tilde{D}\leq 1$.
Let $Q$ denote all the finitely discrete probability measures $Q$ on $\mathcal{X}$, note that
$N(\eta,\tilde{\Pi},L_{2}(\mathbb{P}_{n}))\leq 1+N(\eta,\Pi,L_{2}(\mathbb{P}_{n})),$
and by the result of covering number of VC-class functions (van der Vaart and Wellner, 2023, Theorem 2.6.7), we have
[…]
$\leq$ $CL\sqrt{M\underset{i\leq n}{\max}K_{M}(i)\text{VC}(\Pi)}.$
Taking expectation w.r.t. $\{X_{i},W_{i},\epsilon_{i}\}^{n}_{i=1}$ in (28), by Jensen’s inequality we have
$E\left[\underset{\pi\in\Pi}{\sup}\left|G^{0}_{n}(\pi)\right|\right]\leq E\left[CL\sqrt{M\underset{i\leq n}{\max}K_{M}(i)\text{VC}(\Pi)}\right]\leq CL\sqrt{M\cdot\text{VC}(\Pi)E\left[\underset{i\leq n}{\max}K_{M}(i)\right]}.$ (29)
Since $E[K^{q}_{M}(i)]$ is bounded uniformly in $n$ for any $q>0$ (Abadie and Imbens, 2006, Lemma 3), and $\{K_{M}(i)\}^{n}_{i=1}$ is identically distributed, then by the sharper maximal inequality (Zhang and Chen, 2021, Corollary 7.1), for any $q>1$ we have
$E\left[\underset{i\leq n}{\max}K_{M}(i)\right]=o(n^{\frac{1}{q}}).$ (30)
Therefore, let $c_{6}>0$, there exists $N(c_{6},q)$ such that for all $n>N(c_{6},q)$, we have $E\left[\max_{i\leq n}K_{M}(i)\right]\leq c_{6}n^{\frac{1}{q}}$. Combining (24), (29) and (30), for any $q>1$ and $n>N(c_{6},q)$, we have
$E\left[\underset{\pi\in\Pi}{\sup}\left|\frac{1}{n}\sum^{n}_{i=1}(2W_{i}-1)\left[\sum_{j=1}^{n}\pi(X_{j})M_{ji}\right]\epsilon_{i}\right|\right]\lesssim$ $L\sqrt{\frac{M\cdot\text{VC}(\Pi)}{n^{1-1/q}}},$ (31)
where we use the notation $A\lesssim B$ to denote that $A$ is less than $B$ up to a universal constant.
Step 3.4.2: Concentration inequality
Conditional on $\{X_{i},W_{i}\}_{i=1}^{n}$, let $b_{i}(\pi)=(2W_{i}-1)\left[\sum_{j=1}^{n}\pi(X_{j})M_{ji}\right]$, the term
$\sum\limits_{i=1}^{n}(2W_{i}-1)\left[\sum_{j=1}^{n}\pi(X_{j})M_{ji}\right]\epsilon_{i}=\sum\limits_{i=1}^{n}b_{i}(\pi)\epsilon_{i}$
is the sum of independence and $E\left[b_{i}(\pi)\epsilon_{i}\right]=0$. Note that
$\left|\frac{b_{i}(\pi)\epsilon_{i}}{2L\cdot\underset{i\leq n}{\max}K_{M}(i)}\right|\leq\frac{\sum_{j=1}^{n}\pi(X_{j})M_{ji}}{\underset{i\leq n}{\max}K_{M}(i)}\leq\frac{\sum_{j=1}^{n}M_{ji}}{\underset{i\leq n}{\max}K_{M}(i)}=\frac{K_{M}(i)}{\underset{i\leq n}{\max}K_{M}(i)}\leq 1.$
Given $\{X_{i},W_{i}\}_{i=1}^{n}$, the unit-level terms $\left\{\frac{b_{i}(\pi)\epsilon_{i}}{2L\cdot\max_{i\leq n}K_{M}(i)}\right\}_{i=1}^{n}$ satisfy the conditions of Talagrand’s inequality (Massart, 2007, Page 169-170). Therefore, it follows that for any $t>0$, with at least probability $1-\exp(-t)$, we have
$\underset{\pi\in\Pi}{\sup}\left|\sum_{i=1}^{n}\frac{b_{i}(\pi)\epsilon_{i}}{2L\cdot\underset{i\leq n}{\max}K_{M}(i)}\right|\leq$ $E_{\epsilon}\left[\underset{\pi\in\Pi}{\sup}\left|\sum_{i=1}^{n}\frac{b_{i}(\pi)\epsilon_{i}}{2L\cdot\underset{i\leq n}{\max}K_{M}(i)}\right|\right]+2t$
+2($+2\sqrt{\left(2V+16E_{\epsilon}\left[\underset{\pi\in\Pi}{\sup}\left|\sum_{i=1}^{n}\frac{b_{i}(\pi)\epsilon_{i}}{2L\cdot\underset{i\leq n}{\max}K_{M}(i)}\right|\right]\right)t},$
where
$V=\underset{\pi\in\Pi}{\sup}\sum_{i=1}^{n}E_{\epsilon}\left[\frac{b^{2}_{i}(\pi)\epsilon^{2}_{i}}{4L^{2}\cdot\left[\underset{i\leq n}{\max}K_{M}(i)\right]^{2}}\right]\leq\underset{\pi\in\Pi}{\sup}\frac{\sum_{i=1}^{n}b^{2}_{i}(\pi)}{\left[\underset{i\leq n}{\max}K_{M}(i)\right]^{2}}\leq\frac{\sum_{i=1}^{n}K_{M}^{2}(i)}{\left[\underset{i\leq n}{\max}K_{M}(i)\right]^{2}}$
and $E_{\epsilon}(\cdot)$ denote the expectation of $\epsilon$. Note that for any $\alpha>0$,
2($2\sqrt{\left(2V+16E_{\epsilon}\left[\underset{\pi\in\Pi}{\sup}\left|\sum_{i=1}^{n}\frac{b_{i}(\pi)\epsilon_{i}}{2L\cdot\underset{i\leq n}{\max}K_{M}(i)}\right|\right]\right)t}\leq$ $2\sqrt{2Vt}+8\sqrt{E_{\epsilon}\left[\underset{\pi\in\Pi}{\sup}\left|\sum_{i=1}^{n}\frac{b_{i}(\pi)\epsilon_{i}}{2L\cdot\underset{i\leq n}{\max}K_{M}(i)}\right|\right]t}$
$\leq$ $2\sqrt{2Vt}+\alpha E_{\epsilon}\left[\underset{\pi\in\Pi}{\sup}\left|\sum_{i=1}^{n}\frac{b_{i}(\pi)\epsilon_{i}}{2L\cdot\underset{i\leq n}{\max}K_{M}(i)}\right|\right]+\frac{16t}{\alpha}.$
Without loss of generality, we let $\alpha=1$. It follows that, conditional on $\{X_{i},W_{i}\}^{n}_{i=1}$, with at least probability $1-\exp(-t)$, we have
$\underset{\pi\in\Pi}{\sup}\left|\frac{1}{n}\sum_{i=1}^{n}b_{i}(\pi)\epsilon_{i}\right|\leq 2E_{\epsilon}\left[\underset{\pi\in\Pi}{\sup}\left|\frac{1}{n}\sum_{i=1}^{n}b_{i}(\pi)\epsilon_{i}\right|\right]+\frac{36Lt}{n}\underset{i\leq n}{\max}K_{M}(i)+\frac{4L\sqrt{2t}}{n}\sqrt{\sum_{i=1}^{n}K^{2}_{M}(i)}.$
Since $\{b_{i}(\pi)\}_{i=1}^{n}$ are constants given $\{X_{i},W_{i}\}^{n}_{i=1}$, we replace the $E_{\epsilon}\left[\sup_{\pi\in\Pi}\left|\frac{1}{n}\sum_{i=1}^{n}b_{i}(\pi)\epsilon_{i}\right|\right]$ above by $E\left[\sup_{\pi\in\Pi}\left|\frac{1}{n}\sum_{i=1}^{n}b_{i}(\pi)\epsilon_{i}\right|\right]$. Taking expectation w.r.t. $\{X_{i},W_{i}\}^{n}_{i=1}$, for any $t>0$, with at least probability $1-\exp(-t)$, we have
$\underset{\pi\in\Pi}{\sup}\left|\frac{1}{n}\sum_{i=1}^{n}b_{i}(\pi)\epsilon_{i}\right|\leq 2E\left[\underset{\pi\in\Pi}{\sup}\left|\frac{1}{n}\sum_{i=1}^{n}b_{i}(\pi)\epsilon_{i}\right|\right]+\frac{36Lt}{n}\underset{i\leq n}{\max}K_{M}(i)+\frac{4L\sqrt{2t}}{n}\sqrt{\sum_{i=1}^{n}K^{2}_{M}(i)}$ (32)
Let $\delta=\exp(-t)$, it follows from (31) and (32) that for all $n>N(c_{6},q)$, with at least probability $1-\delta$, we have
$\underset{\pi\in\Pi}{\sup}\left|\frac{1}{n}\sum_{i=1}^{n}b_{i}(\pi)\epsilon_{i}\right|\lesssim L\sqrt{\frac{M\cdot\text{VC}(\Pi)}{n^{1-1/q}}}+\frac{L\log\frac{1}{\delta}}{n}\underset{i\leq n}{\max}K_{M}(i)+\frac{L\sqrt{\log\frac{1}{\delta}}}{n}\sqrt{\sum_{i=1}^{n}K^{2}_{M}(i)}.$ (33)
A.7.4 Step 4: High probability bound for $A(\pi^{*})-A(\hat{\pi})$
Combining (14), (15), (A.7.3), (17), (18), (21) and (33), it follows that for any $\delta>0$, with at least $1-9\delta$ and for all $n>\max\{N(c_{3}),N(c_{4},\delta),N(c_{6},q)\}$, we have
$A(\pi^{*})-A(\hat{\pi})\lesssim$ $\frac{L}{\sqrt{M}}\sqrt{\frac{\text{VC}(\Pi)}{n^{1-1/q}}}+\frac{1}{\sqrt{n}}\left(\sqrt{\log\frac{1}{\delta}}+\sqrt{\text{VC}(\Pi)}\right)\left(\sqrt{V^{*}}+L\right)$ (34)
$+\frac{L\sqrt{\log\frac{1}{\delta}}}{Mn}\sqrt{\sum_{i=1}^{n}K^{2}_{M}(i)}+\frac{L\log\frac{1}{\delta}}{Mn}\underset{i\leq n}{\max}K_{M}(i).$
Note that $\frac{1}{n}\sqrt{\sum_{i=1}^{n}K^{2}_{M}(i)}=\frac{1}{\sqrt{n}}\sqrt{\frac{1}{n}\sum_{i=1}^{n}K^{2}_{M}(i)}=O_{p}\left(\frac{1}{\sqrt{n}}\right)=o_{p}\left(\frac{1}{\sqrt{n^{1-1/q}}}\right)$, where the second equality follows from Markov’s inequality. Let $c_{7}>0$, then for any $\delta>0$, there exists $N(c_{7},\delta)$ such that for all $n>N(c_{7},\delta)$, we have
$\Pr\left(\frac{1}{n}\sqrt{\sum_{i=1}^{n}K^{2}_{M}(i)}\leq\frac{c_{7}}{\sqrt{n^{1-1/q}}}\right)\geq 1-\delta.$ (35)
Similarly, since $\frac{1}{n}\max_{i\leq n}K_{M}(i)=o_{p}\left(\frac{1}{n^{1-1/q}}\right)$ by Markov’s inequality. Let $c_{8}>0$, there exists $N(c_{8},\delta)$ such that for all $n>N(c_{8},\delta)$, we have
$\Pr\left(\frac{1}{n}\underset{i\leq n}{\max}K_{M}(i)\leq\frac{c_{8}}{n^{1-1/q}}\right)\geq 1-\delta.$ (36)
Recall that $A(\pi^{*})-A(\hat{\pi})=2R(\hat{\pi})$. Let $N_{\delta,q}=\max\{N(c_{3}),N(c_{4},\delta),N(c_{6},q),N(c_{7},\delta),N(c_{8},\delta)\}$, it follows from (34), (35), and (36) that, for any $\delta>0$ and any $q>1$, with at least probability $1-11\delta$ and for all $n>N_{\delta,q}$, we have
$R(\hat{\pi})\lesssim\frac{L}{\sqrt{Mn^{1-1/q}}}\left(\sqrt{\text{VC}(\Pi)}+\sqrt{\frac{\log\frac{1}{\delta}}{M}}\right)+\frac{1}{\sqrt{n}}\left(\sqrt{\text{VC}(\Pi)}+\sqrt{\log\frac{1}{\delta}}\right)\left(\sqrt{V^{*}}+L\right)+\frac{L\log\frac{1}{\delta}}{Mn^{1-1/q}}.$
Appendix B Additional Information for Simulation Studies
B.1 Data Generating Process
Table S1 presents detailed information for the data generating processes in simulation studies. Additionally, the proportions of the controlled versus treated units are displayed. For the propensity score model, Scenario 1 is the same as Kang and Schafer (2007), where $l(\cdot)$ takes a linear form relative to its argument and the proportion is balanced. In Scenario 2 we consider a nonlinear form of $l(\cdot)$, while the proportion is still balanced. Scenario 3 is modified from Scenario 1 to introduce imbalanced proportion. We consider a randomized experiment in Scenario 4, where the proportion is extremely imbalanced. Scenario 5 is modified from Wu et al. (2020), with a nonlinear propensity score model and extremely imbalanced proportion.
Table S1: Data generating processes for different scenarios in simulation studies. Outcome variable $Y=m(X)+Wc(X)+Z$, where binary treatment variable $W$ generated by $\Pr(W=1|X)=\exp\left(l(X)\right)/\left[1+\exp\left(l(X)\right)\right]$.
Propensity score model $l(X)$ Controlled versus treated
1. Linear, balanced $-X_{1}+0.5X_{2}-0.25X_{3}-0.1X_{4}$ 50:50
2. Nonlinear, balanced $0.1X_{1}^{3}+0.2X_{2}^{3}+0.3X_{3}$ 50:50
3. Linear, imbalanced $2.1-X_{1}+2X_{2}-0.25X_{3}-0.1X_{4}$ 23:77
4. Constant, imbalanced $\log 9$ 10:90
5. Nonlinear, imbalanced $1+\exp(X_{2})+\sin X_{1}\cos X_{3}$ 12:88
Main effect model $m(X)$
1. Linear $1+2X_{1}-X_{2}+0.5X_{3}-1.5X_{4}$
2. Nonlinear $4\sin(X_{1})+2.5\cos(X_{2})-X_{3}X_{4}$
Contrast function $c(X)$
1. Tree $12I\left\{X_{1}>0,X_{2}>0\right\}-$ 1
2. Non-tree $12I\left\{2X_{2}-\exp(1+X_{1})+2>0\right\}-$ 1
B.2 Simulation Results with Nonlinear Main Effect
Figure S1: Boxplot of empirical value functions in different scenarios, where the main effect model is nonlinear and the contrast function is tree. The global optimal value function is 1.77.
Figure S2: Boxplot of empirical value functions in different scenarios, where the main effect model is nonlinear and the contrast function is non-tree. The global optimal value function is 1.87.
B.3 Comparison with Policy Tree
Figure S3: Boxplot of empirical value functions in different scenarios, where the main effect model is linear and the contrast function is tree.
Figure S4: Boxplot of empirical value functions in different scenarios, where the main effect model is linear and the contrast function is non-tree.
Appendix C Additional Results in Real Data Application
(a) Depth-2 decision tree learned by MB-LASSO-M5
(b) Depth-2 decision tree learned by Policy Tree
Figure S5: Optimal policies estimated by using all the samples in DW dataset. (a) The policy derived from MB-LASSO-M5 depends only on pre-treatment earnings, and 401 (90%) of the samples are predicted to have "exposed" as the optimal choice. (b) The policy derived from Policy Tree depends on both pre-treatment earnings and education, and 290 (65%) of the samples are predicted to have "exposed" as the optimal choice.
(a) Depth-2 decision tree learned by MB-M5
(b) Depth-2 decision tree learned by Policy Tree
Figure S6: Optimal policies estimated by using all the samples in DW-CPS3 dataset. (a) The policy derived from MB-M5 depends on age, education and pre-treatment earnings, and 560 (64%) of the samples are predicted to have "exposed" as the optimal choice. (b) The policy derived from Policy Tree depends on age and pre-treatment earnings, and 557 (64%) of the samples are predicted to have "exposed" as the optimal choice.